### v2

# Markov Decision Processes                                  Under External Temporal Processes

Abstract.
Most reinforcement learning algorithms treat the context under which they operate as a stationary, isolated, and undisturbed environment. However, in real world applications, environments constantly change due to a variety of external events. To address this problem, we study Markov Decision Processes (MDP) under the influence of an external temporal process. We formalize this notion and discuss conditions under which the problem becomes tractable with suitable solutions. We propose a policy iteration algorithm to solve this problem and theoretically analyze its performance. We derive results on the sample complexity of the algorithm and study its dependency on the extent of non-stationarity of the environment. We then conduct experiments to illustrate our results in a classic control environment.
1. Introduction
Reinforcement Learning (RL) provides a framework for sequential decision-making problems, wherein an agent has to make decisions at each time step so that a given objective is achieved. A huge amount of research has been done in this field in recent years, resulting in remarkable progress in various applications like robotics, finance, and even video games (Zhao et al., 2020; Charpentier et al., 2021; Mnih et al., 2015).
Central to RL is the mathematical framework of Markov Decision Processes (MDPs). In an MDP, an agent is assumed to be in an ‘environment’, and at each time step, the agent has to take a decision or an ‘action’. The state of the environment changes stochastically in a Markov manner. It depends only on the current state and the agent’s current action.
While most RL research models the agent’s environment as an MDP, this paper argues that this is a very simplistic model of the real world, as one assumes that the state of a Markov Decision Process might constantly be changing, but the rules of the state transition are still fixed. However, in practical applications, external events may influence the agent’s environment and change its dynamics. For example, consider the problem of portfolio management in finance. An agent has access to a state consisting of various fundamental and technical indicators of a select number of financial instruments and has to make decisions, sequentially, on whether to hold on to an instrument, sell it, or short it, etc. But the way the prices of these instruments move may be affected by external forces, such as government decisions like rate hikes or sudden market movements caused by events occurring in a completely different sector of the economy not being considered by the agent. Such an external event might exert lasting influence over different time scales depending on the kind of event that occurred. A rate cut might affect stock prices for a few months, while a flash crash might cease its effect in a few hours.
As another example, consider a navigation problem in which an agent has control of a robot and needs to travel from one point to another. While the agent can learn to do this task in ideal conditions, the optimal actions taken by the agent might depend on external conditions such as rain, sudden traffic changes, or human movement patterns. So, while an agent can learn to achieve goals in an ideal environment, it cannot be oblivious to external changes that influence the environment dynamics differently.
To tackle this problem, we consider a Markov Decision Process that is perturbed by an external discrete-time temporal process. This external process affects the dynamics of the MDP in a non-Markovian fashion, resulting in a non-stationary problem.
Figure 1. An illustration of the lasting non-stationary effect of perturbations due to exogenous events. The plot shows the prices of two assets following a Geometric Brownian motion. A few exogenous Gaussian shocks to one asset forces it to diverge significantly from the other over the long term, with a markedly different pattern of evolution long after the effect of the exogenous event decays.
Contributions
  1. (1)
    We formulate a non-stationary reinforcement learning problem consisting of an MDP perturbed by an external discrete-time temporal process that affects the dynamics of the MDP in a non-Markovian fashion. We describe sufficient conditions under which such a problem has a well-defined solution.
  2. (2)
    We derive conditions under which a solution can be determined approximately by an agent based on just the current state and a finite history of past events. This is possible when the perturbations caused by events older than $t$ time steps on the MDP transition dynamics and the event process itself are bounded in total variation by $M_{t}$ and $N_{t}$ respectively, and $\sum_{t}M_{t},\sum_{t}N_{t}$ are convergent series. We analyze the trade-off between the performance of the agent and the length of history considered.
  3. (3)
    We propose a policy iteration algorithm to solve this problem and theoretically analyze its behavior. We show that policy improvement occurs in regions of the state space where the Bellman error is not too low based on $M_{t}$ and $N_{t}$. Higher the approximation error due to non-stationarity, smaller the region where the policy is guaranteed to improve.
  4. (4)
    We then analyze the sample complexity of policy evaluation in this setting. For the pathwise Least Squares Temporal Difference algorithm with linear function approximation, we extend existing results in the stationary setting to study the impact of non-stationarity on the expected error of the learned value function. We show that with high probability, the expected error of the value function estimate decomposes into approximation error due to discarding old events, inherent error due to linear function approximation, and the standard error terms due to stochasticity and mixing that are present in the stationary case.
  5. (5)
    We further explain our theoretical results using the discrete-time Hawkes process and conduct experiments in a non-stationary pendulum environment to validate our results.
2. Related Work
While most Reinforcement Learning research considers a stationary Markov Decision Process, there has been some work done on tackling various kinds of non-stationarities that might affect the MDP. Many works consider environments that are piece-wise stationary, i.e., the transition and reward functions remain unchanged except during a select few time steps called change points, during which they can change arbitrarily. Different works consider different assumptions on prior knowledge about the nature of the change points etc., available to the agent.
Padakandla et al. (2020) assume that the transition and reward functions belong to a fixed set of ‘contexts’ between change points. They propose a model-free algorithm that uses a change point detection algorithm that determines when a change occurs by analyzing the state-reward-next state tuples obtained from the environment. However, their algorithm assumes knowledge of the pattern of changes in the context.
Alegre et al. (2021) propose a model-based algorithm that can deal with arbitrarily different types of contexts and continuous state and action spaces. Other works consider drifting environments wherein the transition and reward functions can vary at any time step, while satisfying some constraints to make the problem tractable. Lecarpentier and Rachelson (2019) consider a finite horizon MDP in which the transition dynamics and reward function evolve such that they are Lipschitz continuous with respect to time. Cheung et al. (2020) consider such a setting in a finite MDP where the amount of total change in the environment, referred to as the variation budget, is known to the agent. They propose an optimistic Value Iteration-based algorithm and prove guarantees on the upper bound of its dynamic regret.
Wei and Luo (2021) propose a black-box reduction that turns any suitable algorithm with optimal performance in a near-stationary environment into an algorithm that achieves optimal dynamic regret in a non-stationary environment.
Hadoux et al. (2014) formalized a class of piecewise stationary MDPs called Hidden-Semi-Markov-Model MDPs where the transition from one hidden stationary MDP mode to another is semi-Markov in nature. To solve this problem, they adapt the Partially Observable Monte Carlo Planning (POMCP) algorithm used for solving Partially Observable Markov Decision Processes (POMDPs).
Some algorithms assume additional structure on the nature of non-stationarity. Feng et al. (2022) propose Factored Non-stationary MDPs (FN-MDPs), where the transition dynamics are defined in terms of a Dynamic Bayesian Network over the states, actions, rewards, and latent change factors that cause non-stationarity. Here, both the states and the latent change factors have transition dynamics obeying this causal graph. They propose using a Variational Auto-Encoder (VAE) framework to learn transition dynamics and present results in MuJoCo and robotic manipulation tasks.
Another factorization setting with a resemblance to ours has been proposed by Efroni et al. (2022) called the Exogenous MDP, wherein the environment state is divided into two parts, an endogenous state which is controllable by the agent’s actions, and an exogenous state, which evolves on its own and does not affect the agent’s reward. The precise way in which this division exists is unknown to the agent. While this may seem similar to our formulation of states and external events, in our case, the external events do influence the dynamics of the state and thereby, the rewards obtained by the agent.
A similar decomposition of the state space into endogenous and exogenous sub-spaces was studied by Dietterich et al. (2018), who assumed that the exogenous states evolve in a Markov fashion and the reward function can be additively decomposed into an exogenous reward and an endogenous reward. While their work studies how to decompose the given state into endogenous and exogenous components, in contrast, our work deals with essentially the opposite task, which is to study the effect of incorporating exogenous event information into reinforcement learning algorithms.
3. Preliminaries
A Markov Decision Process (MDP) is defined by a tuple $(S,A,\left\{A(s):s\in S\right\},Q,c,\gamma)$, where $S$ is the state space, the set of all possible states of the environment, $A$ is the action space, the set of all possible actions that can be taken by the agent, $A(s)$ is the set of allowed actions at state $s$, $Q$ is the transition probability kernel, $c$ is the cost function, and $\gamma$ is the discount factor.
At each time step $t$, the agent is in state $s\in S$ and has to take action $a\in A(s)\subseteq A$. Depending on the action taken, a cost $c(s,a)$ is incurred by the agent, and the environment transitions to state $s^{\prime}\in S$ sampled from $Q\left(.|(s,a)\right)$, which is then communicated to the agent. This process repeats indefinitely.
In general, $S$ and $A$ are Borel spaces, $A(s)$ is a non-empty measurable subset of $A$ for each $s\in S$, $c$ is a measurable function, and $\gamma\in(0,1)$.
$Q$ is a stochastic kernel on $S$ given $\mathbb{K}$, where $\mathbb{K}=\left\{(s,a):s\in S,a\in A(s)\right\}$ is the set of all feasible state-action pairs. That is, $Q(.|y)$ is a probability measure on $S$ for each $y\in\mathbb{K}$, and $Q(B|.)$ is a measurable function on $\mathbb{K}$ for each $B\in\mathcal{B}(S)$, the Borel $\sigma$-algebra on $S$.
Any rule $\pi=\{\pi_{t}\}$ by which the agent chooses the action, as a function of state and possibly time, is called a policy. A policy can be either deterministic with $\pi_{t}(s)=a_{t}\in A(s)$, or stochastic with $\pi_{t}(.|s\>)$ being a stochastic kernel on $A(s)$ given $s\in S$.
A policy that is independent of time is called a stationary policy. The goal of the agent is to determine a policy $\pi$ among all possible policies $\Pi$ such that the following performance criterion is minimized:
$V^{\pi}(s)=\mathsf{E}_{s_{t},a_{t}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],\qquad\pi\in\Pi,x\in S,$
the sum of discounted costs incurred by the agent when it starts from a state and follows policy $\pi$ at each time step, resulting in a trajectory of states $s_{0}=s,s_{1},s_{2},\dots,$ by taking actions $a_{0},a_{1},a_{2},\dots$. The function $V^{\pi}$ is called the value function of the policy $\pi$.
By the Monotone Convergence Theorem, the above can be written as the limit
$V^{\pi}(s)=\lim_{n\rightarrow\infty}V_{n}^{\pi}(s),$
where $V_{n}$ is the $n$-stage cost, the sum of discounted costs obtained from the first $n$ time steps.
A policy $\pi^{*}$ that achieves the minimum value is called an optimal policy and the corresponding value function $V^{\pi^{*}}$ is denoted as $V^{*}$.
3.1. Bellman optimality equation
A measurable function $v:S\rightarrow\mathbb{R}$ is said to be a solution to the Bellman optimality equation if it satisfies:
$v(s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
The right-hand side of this equation can be considered as an operator $\mathcal{T}$ acting on the function $v$, leading to the optimal Bellman equation
$v=\mathcal{T}v,$
where the optimal Bellman operator $\mathcal{T}$ is defined as
$[\mathcal{T}v](s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
A function that satisfies the optimal Bellman equation is a fixed point of the operator $\mathcal{T}$. Under suitable conditions, the optimal value function satisfies this equation.
Assumption 1.
  1. (1)
    The one-stage cost $c$ is lower semi-continuous, non-negative and inf-compact on $\mathbb{K}$.
  2. (2)
    Q is strongly continuous.
Assumption 2.
There exists a policy $\pi$ such that $V^{\pi}(s)<\infty$ for all $s\in S$.
Under the above assumptions, we can make some statements about the existence and uniqueness of the optimal policy and value function, given by the following result.
Theorem 3.1 (Theorem 4.2.3 of Hernández-Lerma and Lasserre (2012)).
Suppose the above two assumptions hold, then:
  1. (1)
    The optimal value function $V^{*}$ is the point-wise minimal solution to the above Bellman equation. If $u(.)$ is any other solution to the above equation, then $u(.)\geq V^{*}(.)$.
  2. (2)
    There exists a function $\pi^{*}:S\rightarrow A$ such that $\pi^{*}(s)\in A(s)$ that attains the minimum in the Bellman equation, i.e.,
    $V^{*}(s)=c(x,\pi^{*}(s))+\gamma\int V^{*}(s^{\prime})Q(ds^{\prime}|s,\pi^{*}(s))\quad\text{for all }s\in S,$
    and the deterministic stationary policy $\pi^{*}$ is optimal. Conversely, any deterministic stationary policy that is optimal satisfies the above equation.
  3. (3)
    If an optimal policy exists, then there exists one that is deterministic and stationary.
4. MDP under an external temporal process
Consider an MDP $\mathcal{M}=(S,A,Q,r,\gamma)$ where
    $S$: The state space, a closed subset of $\mathbb{R}^{m}$
    $A$: The action space, a closed subset of $\mathbb{R}^{n}$
    $\mathbb{K}=S\times A$: The set of all feasible state-action pairs
    $r:S\times A\times S\rightarrow\mathbb{R}$: Reward function as a function of the state-action pair and the resultant next state
    $\gamma$: Discount factor
    $Q$: A stochastic kernel on $S$ given $\mathbb{K}$
We assume that all actions in $A$ are allowed at every state in $S$, so there is no need to specify the allowed action set separately. Also, instead of a cost function $c$, we consider a reward function $r$, where the problem objective is maximize the discounted sum of rewards instead.
Consider an external discrete-time temporal process that is independent of the above MDP but affects the transition probabilities of the MDP.
The probability of an event occurring in the external process at time $t$ is a function of all the previous (external) events till that time. Associated with each event is a probabilistic mark $x\in\mathcal{X}$, where $\mathcal{X}$ is a closed subset of $\mathbb{R}^{d}$.
Suppose at time $t$ that the state of the MDP is $s\in S$. Suppose the history of external events till that time is $H_{t}=\{(t_{1},x_{1}),(t_{2},x_{2}),\dots,(t_{j},x_{j})\}$, where the ordered pair $(t_{i},x_{i})$ denotes an event with mark $x_{i}$ that occurred at time $t_{i}$.
Then, this history perturbs the stochastic kernel $Q$ of the MDP to give rise to a new transition kernel $Q_{H_{t}}$ on $S$ given $\mathbb{K}$.
Similarly, this history also determines the next event that occurs in the temporal process. We denote the distribution of the event mark at time $t$ by $Q^{X}_{H_{t}}$, which is a probability distribution on $\mathcal{X}$. Since this describes the event distribution that is external to the MDP, it does not depend on the state of the MDP or the action taken by the agent.
4.1. Assumptions
4.1.1. On the external process
  1. (i)
    There exists an event with mark, say $X=0$ (zero), that is equivalent to a non-event.
  2. (ii)
    Old events have a reduced influence on the current probability transition kernel of the MDP. More specifically, there exists a convergent series $\sum_{T}M_{T}$ of real numbers, such that at any time $t$, for any event histories $H_{t}$ and $H^{\prime}_{t}$, if the MDP is in state $s$ and action $a$ is taken, and if $H_{t}$ and $H^{\prime}_{t}$ differ by only one event at time $t^{\prime}$, then
    $\mathsf{TV}\left(Q_{H_{t}}(.|s,a),Q_{H^{\prime}_{t}}(.|s,a)\right)<M_{T},\quad\text{for all }(s,a)\in\mathbb{K},$
    provided $t-t^{\prime}\geq T$. That is, if an event is older than $T$, its influence on the current probability distribution has a Total Variation distance upper bounded by $M_{T}$.
  3. (iii)
    Similarly, old events also have a reduced influence on the probability distribution of the current event mark. Specifically, there exists a convergent series $\sum_{T}N_{T}$ of real numbers such that for event histories $H_{t}$ and $H^{\prime}_{t}$ at time $t$ that differ at only one event at time $t^{\prime}$, if $t-t^{\prime}\geq T$, then
    $\mathsf{TV}\left(Q^{X}_{H_{t}},Q^{X}_{H^{\prime}_{t}}\right)<N_{T}.$
4.1.2. Regularity
Define a new function $r_{P}:\mathbb{K}\rightarrow\mathbb{R}$ induced from $r$ by $P$ as
$r_{P}(s,a)=\mathsf{E}_{s^{\prime}\sim P(.|s,a)}\left[r(s,a,s^{\prime})\right],$
where $P$ is any stochastic kernel on $S$ given $\mathbb{K}$. We assume the following regularity conditions on $r$ and $Q$:
  1. (i)
    $r$ is a bounded measurable function. Without any loss in generality, assume its co-domain is $[0,1]\subset\mathbb{R}$.
  2. (ii)
    The function $r_{P}(s,.)$ is upper semi-continuous for each state $s\in S$ and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process.
  3. (iii)
    $r_{P}$ is sup-compact on $\mathbb{K}$, i.e., for every $s\in S,r\in\mathbb{R}$, and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process, the set $\{a\in A:r_{P}(s,a)\geq r\}(\subseteq A)$ is compact.
  4. (iv)
    $Q$ is strongly continuous.
  5. (v)
    $Q_{H}$ and $Q^{X}_{H}$ are strongly continuous for any feasible event history $H$.
Note: A transition kernel $P$ is said to be strongly continuous if
$v(s,a)=\mathsf{E}_{s^{\prime}\sim P(.|s,a)}[v(s^{\prime})]$
is continuous and bounded on $\mathbb{K}$ for every measurable bounded function $v$ on $S$.
4.2. Examples
Non-Markov self-exciting event processes that can be described as above are studied extensively in finance, economics, epidemiology, etc., in the form of variants of Hawkes and jump processes. While they are studied predominantly in continuous time and sometimes in discrete time, we consider discrete time versions to fit into the traditional discrete-time reinforcement learning framework.
Consider a discrete-time marked Hawkes process $\left(E_{t},X_{t}\right)_{n\in\mathbb{N}}$ defined as follows:
$\left(E_{t}\right)_{t\in\mathbb{N}}$ is a sequence of random variables taking values in $\left\{0,1\right\}$, with $E_{t}$ sampled from $\textsf{Bernoulli}\left(p_{t}\right)$, where the intensity $p_{t}$ at time $t$ depends on the realizations of $E_{t}$ at the previous time steps as:
$p_{1}$ $=\alpha_{0},$
$p_{t}$ $=\alpha_{0}+\sum_{t^{\prime}=1}^{t-1}\alpha_{t-t^{\prime}}E_{t^{\prime}}\;\text{ for }t>1,$
and the marks $X_{t}$ are real-valued Gaussian random variables satisfying
$X_{1}$ $\sim\mathcal{N}(0,1),$
$X_{t}$$\begin{dcases}\sim\mathcal{N}\left(\sum_{t^{\prime}=1}^{t-1}\beta_{t-t^{\prime}}E_{t}X_{t},\;1\right)&\text{ if }E_{t}=1,\\ =0&\text{ otherwise,}\end{dcases}\text{ for }t>1,$
where $(\alpha_{t})$ and $(\beta_{t})$ are sequences that converge sufficiently quickly to zero.
The sum $S_{t}=\sum_{t^{\prime}=1}^{t}E_{t}$ is a self-exciting counting process that is the discrete-time counterpart of the continuous-time Hawkes process.
The above defined sequence of random variables satisfies Assumptions (i) and (ii) required for the external process in Section 4.1.1. $X_{t}$ is zero when no event occurs, and old events have a reduced influence on current events. More precisely, given two histories $(X_{t^{\prime}})_{t^{\prime}<t}$ and $(X^{\prime}_{t^{\prime}})_{t^{\prime}<t}$ that only differ by one event at some time $t^{\prime}\leq t-T$ with $E^{\prime}_{t^{\prime}}=X^{\prime}_{t^{\prime}}=0$ and $E_{t^{\prime}}=1$, $X_{t^{\prime}}=x$, the difference in intensities is $p_{t}-p^{\prime}_{t}=\alpha_{t-t^{\prime}}$, and so,
[…]
$=\alpha_{t-t^{\prime}}+p_{t}\;\textsf{erf}\left(\frac{\beta_{t-t^{\prime}}}{2\sqrt{2}}\right)$
$<\alpha_{T}+\textsf{erf}\left(\frac{\beta_{T}}{2\sqrt{2}}\right)=N_{T}.$
For sufficiently quickly converging $\alpha_{T}$ and $\beta_{T}$, the total variation perturbation $N_{T}$ due to events older than $T$ time steps goes to zero, satisfying Assumption (ii).
While $(\alpha_{n})$ and $(\beta_{n})$ are any general sequences that need to satisfy some regularity conditions (Seol, 2015) and the above bound, they could also decay in a more structured manner. For instance, for some $c>0.5,\lambda>0$ letting
$\alpha_{n}=ce^{-\lambda n},$
gives a process that has exponentially decaying intensity due to past events. In the terminology employed to describe the standard continuous-time Hawkes process, $\alpha_{0}=c$ is like the base or background intensity, and $e^{-\lambda x}$ is the excitation function. Similarly, $\beta_{n}$, which determines the distribution of the event, could also be a parametric sequence that decays in an exponential or polynomial fashion.
5. Characterization of the problem setting
Given the MDP $(S,A,r,\gamma,Q)$ and an external discrete-time temporal process $\{(t_{i},X_{i})\}_{i}$ that acts on this MDP, we construct a new augmented MDP where the marks of all the events that occurred till time $t$ are appended to the state $s\in S$ to form a new augmented state $\bar{s}\in\bar{S}$, where
$\bar{s}=(s,x_{t},x_{t-1},x_{t-2},x_{t-3},\dots),\qquad\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}\left(\subset\mathbb{R}^{m}\times\prod_{t=0}^{\infty}\mathbb{R}^{d}\right).$
Here, $x_{t}$ is the mark of the external event that occurred at time $t$. If no event occurred at time $t$, then $x_{t}=0$.
Because of the interference from external events, the decision process $(S,A,r,\gamma,Q)$ does not remain an MDP since the transition function is being perturbed by external events. However, the corresponding decision process $\mathcal{M}_{X}=(\bar{S},A,r,\gamma,Q_{H})$ is a Markov Decision Process because all the factors that affect the transitions are included in the augmented state $\bar{s}\in\bar{S}$.
Now, the state space of this MDP is infinite-dimensional and poses several challenges. However, since we assume that very old events do not affect the transition probabilities during the current time step, we should be able to find a suitable policy that only depends on a finite horizon of past events. This is formalized by the following result:
Theorem 5.1.
Suppose the MDP and the external process satisfy the assumptions in 4.1.1 and 4.1.2.
  1. (1)
    There exists a deterministic optimal policy for the augmented MDP and a corresponding optimal value function $V^{*}$ that satisfies the optimal Bellman equation.
  2. (2)
    For any $\epsilon>0$, there exists a horizon $T$ and a policy $\pi^{(T)}$ that depends only on the current state and past $T$ events such that
    $V^{\pi^{(T)}}\geq V^{*}-\epsilon.$
    That is, for every required maximum suboptimality $\epsilon$, there exists a corresponding “finite-horizon” policy that achieves that $\epsilon$-suboptimal value.
    The magnitude of $T$ depends on the rate of convergence of the series $\sum_{t}M_{t}$ and $\sum_{t}N_{t}$.
To prove this, we use a Lemma about the properties of a class of auxiliary MDPs $\mathcal{M}^{(T)}$, whose proof is provided in the Appendix.
For each time $T>0$, we construct MDP $\mathcal{M}^{(T)}$ as having the same underlying stochastic process as $\mathcal{M}_{X}$, with only one difference: only events in the past $T$ time steps affect the current state transition, and all the events occurring before that have no effect and can effectively be replaced by zeroes. Depending on the context, the state space of this MDP can either be taken as $S\times\prod_{t=0}^{\infty}\mathcal{X}$ or $S\times\prod_{t=0}^{T}\mathcal{X}$, since all the older events are zeroed out.
Any policy $\pi$ for this MDP has an equivalent policy in $\mathcal{M}_{X}$ and vice-versa, where any extra elements in the state are ignored or considered zero.
Lemma 5.2.
Let $\pi$ be an arbitrary policy as a function of the augmented state $\bar{s}\in\bar{S}$. That is, $\pi$ can be either stochastic or deterministic, and can depend on the current state $s\in S$ and any number of past events. Then,
$\left|\left|V^{\pi}_{\mathcal{M}^{(T)}}-V^{\pi}_{\mathcal{M}_{X}}\right|\right|_{\infty}\leq\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{\pi}_{\mathcal{M}^{(T)}}||_{\infty}\right)\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
where $V^{\pi}_{\mathcal{M}}$ denotes the value function of policy $\pi$ in MDP $\mathcal{M}$.
5.1. Proof of Theorem 5.1
Part (1) follows from the assumptions and Theorem 3.1. For part (2), let $\pi^{*}$ be the deterministic optimal stationary policy for $\mathcal{M}_{X}$ and let $\pi^{*(T)}$ be the deterministic optimal stationary policy for $\mathcal{M}^{(T)}$. From Lemma 5.2, since $||r||_{\infty}\leq 1$ and $||V||_{\infty}\leq\frac{1}{1-\gamma}$,
$V^{\pi^{*(T)}}$ $\geq V^{\pi^{*(T)}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\geq V^{\pi^{*}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\geq V^{\pi^{*}}_{\mathcal{M}_{X}}-\frac{2}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
Since both series $\sum_{t}M_{t}$ and $\sum_{t}N_{t}$ converge, there exists $T\in\mathbb{N}$ such that
$\sum_{t=T+1}^{\infty}M_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}\text{ and}\sum_{t=T+1}^{\infty}N_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}.$
Choosing such a $T$ gives the Theorem.
6. A policy iteration Algorithm
Now that an $\epsilon$-optimal policy exists that is a function of only a finite event horizon, we wish to find such a policy. We propose a policy iteration algorithm, which alternates between approximate policy evaluation and approximate policy improvement. The overall algorithm is given in 1.
Start with arbitrary deterministic policy $\pi_{0}:S\times\mathcal{X}^{T+1}\rightarrow A$;
foreach $k\in\{0,1,\dots\}$ till termination do
       // Approximate Policy Evaluation
$\hat{V}_{k}\left((s,x_{0:\infty})\right)=\mathsf{E}_{x^{\prime}_{T+1:\infty}\sim\mu_{1}}V^{\pi_{k}}\left((s,x_{0:T},x^{\prime}_{T+1:\infty})\right).$
      // Approximate Policy Improvement
$\pi_{k+1}\left((s,x_{0:\infty})\right)=\mathop{\mathrm{argmax}}\limits_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\bigg{[}r(s,a,s^{\prime})+\gamma\hat{V}_{k}\left((s^{\prime},x^{\prime},x_{0:T-1})\right)\bigg{]}.$
Algorithm 1 Policy Iteration
For the policy evaluation step, we only wish to consider candidate value functions that are functions of the finite event horizon, which is why it is not an exact step. In this approximate policy evaluation step, we define the value at an infinitely augmented state as a function of the state augmented by its finite event horizon, by sampling the previous older events from some distribution $\mu_{1}$. This $\mu_{1}$ is theoretically arbitrary and empirically can be based on the actual event process. The value function can thus be approximated using Monte Carlo methods. For the purpose of analysis, we consider the exact evaluation of the approximate function.
Similarly, in the policy improvement step, the policy is improved based on the reward and value function resulting from one transition, and this transition can be due to past events coming from any arbitrary distribution $\mu_{2}$. The deterministic policy at an augmented state depends only on the finite event horizon and is just the action that maximizes the RHS that depends on the approximate value function.
Solving the optimization problem in the policy improvement step requires knowledge of the models $r(s,a,s^{\prime})$, $Q$ and $Q^{X}$. In this work, we ignore any approximation errors arising due to estimating the models and analyze the algorithm assuming the exact step is taken as defined.
6.1. Analysis
For the analysis of this algorithm, we need the following Lemma, which bounds the difference in the value of the value function at two different augmented states when they differ in events that are more than $T$ time steps old.
Lemma 6.1.
Let $\pi$ be a policy that depends only on the current state and the past $T$ events. Then
$\sup_{\bar{s}=(s,x_{0:\infty})}\left|V^{\pi}(\bar{s})-V^{\pi}((s,x_{0:T},(0)_{T+1:\infty}))\right|<\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
Corollary 6.2.
For a given policy $\pi$ that is a function of events only in the past $T$ time steps, and for any two augmented states $\bar{s}_{1}$ and $\bar{s}_{2}$ that differ in any number of events that occurred before the previous $T$ time steps,
$\left|V^{\pi}(\bar{s}_{1})-V^{\pi}(\bar{s}_{2})\right|\leq\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$ (1)
Throughout the rest of this section, we let $\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)$, the change in the value function induced by changing an arbitrary number of sufficiently old events.
Now, we characterize the behavior of the policy iteration procedure described previously. In general, exact policy iteration results in a sequence of policies that converge to the optimal policy. However, our approximate policy iteration produces policies whose value functions satisfy the following Lemma.
Lemma 6.3.
$V^{\pi_{k+1}}\geq V^{\pi_{k}}-\frac{2+5\gamma}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
That is, there is no guarantee that the sequence of value functions is non-decreasing. The only guarantee is that they do not decrease by more than the specific amount given by the Lemma. This is due to the approximation error $\epsilon$ at each time step.
In general, approximate policy iteration algorithms converge under the condition that the approximation errors slowly vanish over the course of the algorithm. However, our approximation error of $\epsilon$ remains constant throughout, leading to the above situation.
However, intuitively it would seem that such a small approximation error $\epsilon$ should only disturb the convergence of the algorithm when we are $\epsilon$-close to the optimal solution. During the initial iterations when we are very far from the solution, such small approximation errors should not matter.
It turns out that this is actually the case in some sense. Degradation of the value function occurs only in parts of the state space where the Bellman error is close to zero. A Bellman error of zero generally, although not always, corresponds to the optimal policy. This interplay between Bellman error and approximation error in our algorithm is formalized by the following result.
Theorem 6.4.
At iteration $k$, for any augmented state $\bar{s}\in S\times\prod_{t=0}^{\infty}\mathcal{X}$, at least one of the following holds:
  1. (1)
    $V^{\pi_{k+1}}(\bar{s})\geq V^{\pi_{k}}(\bar{s})$.
  2. (2)
    $|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})|<\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$.
That is, the performance of the policy improves everywhere except the region of the state space where the Bellman error is very small. The faster the decay of exogenous event influence, the lesser the approximation error and smaller the region of state space where the policy may not improve.
Proof.
Which of the above statements holds depends on whether or not the state $\bar{s}$ falls in the “sub-optimality” set $O$, which is the set of all states $\bar{s}=(s,x_{0:\infty})\in\bar{S}$ that satisfy
$\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]$
$<\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]-\delta$
for some suitable $\delta>0$ yet to be chosen. It is to be noted that whether or not this condition holds, the above inequality or the corresponding equality holds for $\delta=0$ by the definition of $\pi_{k+1}$.
Suppose the above condition holds. Then,
$V^{\pi_{k}}((s,x_{0:\infty}))$ $\leq\mathsf{E}_{x^{\prime}_{T+1:\infty}\sim\mu_{2}}V^{\pi_{k}}((s,x_{0:T},x^{\prime}_{T+1:\infty}))+\epsilon$
$=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma V^{\pi_{k}}((s^{\prime},x^{\prime},x_{0:T},x^{\prime}_{T+1:\infty}))\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\epsilon$
$\leq\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\gamma\epsilon+\epsilon$
$<\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k+1}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\gamma\epsilon+\epsilon-\delta$
$\leq\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:\infty}}(.)\\ a=\pi_{k+1}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]$
$\qquad\qquad+\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\epsilon-\delta$
$=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:\infty}}(.)\\ a=\pi_{k+1}(\bar{s})\end{subarray}}r(s,a,s^{\prime})+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:\infty}}(.)\\ a=\pi_{k+1}(\bar{s})\end{subarray}}V^{\pi_{k}}((s^{\prime},x^{\prime},x_{0:\infty}))$
$\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\gamma\epsilon+\epsilon-\delta$
That is, $V^{\pi_{k}}(\bar{s})$ can be bounded recursively in terms of the reward obtained using $\pi_{k+1}$ and $V^{\pi_{k}}(\bar{s}^{\prime})$, where $\bar{s}^{\prime}$ comes from following policy $\pi_{k+1}$. This inequality can further be unrolled infinitely to obtain
$V^{\pi_{k}}(\bar{s})$ $\leq\mathsf{E}_{\pi_{k+1}}\sum_{t=0}^{\infty}\gamma^{t}r(s,a,s^{\prime})-\delta$
$\qquad\qquad+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
$=V^{\pi_{k+1}}(\bar{s}),$
for $\delta=\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\frac{(1+2\gamma)}{1-\gamma}\epsilon$.
Now, suppose $\bar{s}\notin O$. We need to show that this implies condition (2) in the theorem statement, that is, the Bellman error should be very small. It is anyway always true that $\mathcal{T}V\geq V$. In the other direction, for any $\bar{s}=(s,x_{0:\infty})\in\bar{S}$,
$\mathcal{T}V^{\pi_{k}}(\bar{s})$ $=\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:\infty}}(.)\end{subarray}}\left[r(s,a,s^{\prime})+\gamma V^{\pi_{k}}(s^{\prime},x^{\prime},x_{0:\infty})\right]$
$\leq\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:\infty}}(.)\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}(s^{\prime},x^{\prime},x_{0:T-1})\right]+\gamma\epsilon$
$\leq\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}(s^{\prime},x^{\prime},x_{0:T-1})\right]+\gamma\epsilon$
$\qquad\qquad\qquad\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
$=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k+1}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}(s^{\prime},x^{\prime},x_{0:T-1})\right]+\gamma\epsilon$
$\qquad\qquad\qquad\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
$\leq\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}(s^{\prime},x^{\prime},x_{0:T-1})\right]+\gamma\epsilon$
$\qquad\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\delta\qquad(\text{since }\bar{s}\notin O)$
$=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ x^{\prime\prime}_{T:\infty}\sim\mu_{1}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma V^{\pi_{k}}(s^{\prime},x^{\prime},x_{0:T-1},x^{\prime\prime}_{T:\infty})\right]+\gamma\epsilon$
$\qquad\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\delta$
$\leq\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:\infty}}(.)\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma V^{\pi_{k}}(s^{\prime},x^{\prime},x_{0:\infty})\right]+2\gamma\epsilon$
$\qquad\qquad\qquad+\frac{2}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\delta$
$=V^{\pi_{k}}(\bar{s})+\delta+2\gamma\epsilon+\frac{2}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
The first inequality above is due to the definition of $\hat{V}_{k}$, and the second equality is due to the definition of $\pi_{k+1}$. Therefore,
$\left|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})\right|$ $\leq\delta+2\gamma\epsilon+\frac{2}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
$=\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\frac{(1+2\gamma)}{1-\gamma}\epsilon$
$\qquad\qquad\qquad+2\gamma\epsilon+\frac{2}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
$=\sum_{t=T+1}^{\infty}(M_{t}+N_{t})\left[\frac{3-2\gamma+4\gamma}{(1-\gamma)^{2}}+\frac{2(1+2\gamma)}{(1-\gamma)^{3}}\right]$
$=\sum_{t=T+1}^{\infty}(M_{t}+N_{t})\left[\frac{3+2\gamma}{(1-\gamma)^{2}}+\frac{2(1+2\gamma)}{(1-\gamma)^{3}}\right]$
$\leq\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
Proof of Lemma 6.3.
Letting $\delta=0$ in the first part of the above proof gives us the Lemma for which there is no assumption on the extent of policy improvement.
$V^{\pi_{k}}(\bar{s})$ $\leq V^{\pi_{k+1}}(\bar{s})+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
$=V^{\pi_{k+1}}(\bar{s})+\left[\frac{1}{(1-\gamma)^{2}}+\frac{2(1+2\gamma)}{(1-\gamma)^{2}}\right]\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
$=V^{\pi_{k+1}}(\bar{s})+\frac{2+5\gamma}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
In Theorem 6.4, in regions of the augmented state space without guaranteed policy improvement, the Bellman error depends on the approximation error due to ignoring old events, and is a cumulative effect of all events older than $T$ on the current events as well as state transition kernel. Specifically, it is proportional to
$\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
where $T$ is the time window beyond which events are discarded, $M_{t}$ is an upper bound on the total variation disturbance on the state transition kernel caused by an event older that $t$ time steps, and similarly $N_{t}$ bounds the total variation effect due to old events on the event process itself. The faster the decay of exogenous event influence, the smaller the truncation error due to non-stationarity.
For example, consider a discrete-time Hawkes process with an exponentially decaying excitation function $c_{\alpha}e^{-\lambda_{\alpha}t}$ as described in Section 4.2, along with an exponentially decaying $\beta_{t}=c_{\beta}e^{-\lambda_{\beta}t}$ and state transition perturbation that decays as $M_{t}=\bar{c}e^{-\bar{\lambda}t}$. Then, the extra error introduced is proportional to
$\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)$ $\leq\sum_{t=T+1}^{\infty}\left(\bar{c}e^{-\bar{\lambda}t}+c_{\alpha}e^{-\lambda_{\alpha}t}+\textsf{erf}\left(\frac{c_{\beta}e^{-\lambda_{\beta}t}}{2\sqrt{2}}\right)\right)$
$\leq\sum_{t=T+1}^{\infty}\left(\bar{c}e^{-\bar{\lambda}t}+c_{\alpha}e^{-\lambda_{\alpha}t}+\sqrt{1-\exp\left(\frac{-c_{\beta}^{2}}{2\pi}e^{-2\lambda_{\beta}t}\right)}\right)$
$\leq\sum_{t=T+1}^{\infty}\left(\bar{c}e^{-\bar{\lambda}t}+c_{\alpha}e^{-\lambda_{\alpha}t}+\frac{c_{\beta}}{\sqrt{2\pi}}e^{-\lambda_{\beta}t}\right)$
$\leq\int_{t=T}^{\infty}\left(\bar{c}e^{-\bar{\lambda}t}+c_{\alpha}e^{-\lambda_{\alpha}t}+\frac{c_{\beta}}{\sqrt{2\pi}}e^{-\lambda_{\beta}t}\right)\text{d}t$
$=\frac{\bar{c}}{\lambda}e^{-\bar{\lambda}T}+\frac{c_{\alpha}}{\lambda_{\alpha}}e^{-\lambda_{\alpha}T}+\frac{c_{\beta}}{\sqrt{2\pi}\lambda_{\beta}}e^{-\lambda_{\beta}T}.$
From this, it can be seen that as the time window $T$ considered increases, the error reduces as $e^{-\left\{\bar{\lambda},\lambda_{\alpha},\lambda_{\beta}\right\}T}$. So, depending on the rate of influence decay, increasing the time window beyond a certain point gives diminishing returns. The error also has a proportional dependence on $c$, the base intensity of the Hawkes process and an inverse dependency on $\lambda_{\alpha}$, the decay of the excitation function of the Hawkes process.
7. Sample Complexity
The algorithm given in Section 6 defines the value function and its update. But in practice, the value function has to be learnt using samples obtained from the environment. Further, the actual value function is generally approximated from a chosen set of functions.
In this section, we analyze the sample complexity of policy evaluation by using samples from a single sample path induced by the policy, and give error bounds on the evaluation error for the case of a linear function space and some regularity assumptions on the MDP.
Function approximation
Consider a function class $\mathcal{F}$ of functions that are to be used to approximate the value function on the augmented MDP:
$\mathcal{F}\subset\left\{f:\bar{S}\rightarrow\mathbb{R}\right\}.$
Since the domain of the functions in this class is infinite-dimensional, this can be further broken down into two approximations for dealing with these functions practically: a truncation operation to make the domain finite dimensional, and then another class of functions that are then used to cover possible functions on this domain, i.e,
=$\mathcal{F}=\left\{f\;o\;f^{T}_{\text{trunc}}:f\in\mathcal{F}^{T}\right\},$
where $f_{\text{trunc}}:\bar{S}\rightarrow S\times\bigcup_{t=0}^{T}\mathcal{X}$ removes extra events that are older than $T$ time steps while preserving the current state and new events, and $\mathcal{F}^{T}$ is some function class of functions defined on this new truncated domain. For a fixed $f^{T}_{\text{trunc}}$, finding the best function in $\mathcal{F}$ is the same as finding the best function in $\mathcal{F}^{T}$.
Finally, we wish to break down the error between the true value function and the learned truncated value function in terms of the errors introduced due to these two operations as well as the stochasticity of the samples used to learn the approximate function. We wish to bound
$\left|\left|V-\tilde{V}\right|\right|_{\rho},$
where $V$ is the true value function on $\bar{S}$, $\tilde{V}$ is the learned value function in $\mathcal{F}$, and $\rho$ is some distribution over $\bar{S}$, and the norm is the $l^{2}(\rho)$-norm, which is the expected $l^{2}$ norm of the value of the function w.r.t the measure $\rho$, i.e,
$\left|\left|f\right|\right|^{2}_{\rho}=\int_{\bar{S}}f(x)^{2}\;\rho(\mathrm{d}x).$
While the $f^{T}_{trunc}$ operator just ignores old events, we also define a projection operator $\Pi^{trunc}$ onto the function space defined on the truncated domain, as
$\Pi^{trunc}V=\inf_{f:S\times\bigcup_{t=0}^{T}\mathcal{X}}\left|\left|f-V\right|\right|_{\rho}.$
which finds the best approximation that does not depend on the old events, where ‘best’ is defined in expectation with respect to $\rho$.
Linear Function Approximation
A standard approximating function space considered is a an arbitrary finite dimensional function space. Specifically, assume that there are $d$ basis functions $\varphi_{i}:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}$ that linearly span $\mathcal{F}$ and are bounded by $L$, that is, for each $f\in\mathcal{F}$, there exist $\alpha\in\mathbb{R}^{d}$ such that $f=\sum_{i\in[d]}\alpha_{i}\varphi_{i}$. These basis functions together form a feature representation function $\phi:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}^{d}$, defined by $\phi(x)=(\varphi_{1}(x),\dots,\varphi_{d}(x))$.
7.1. Approximate policy evaluation
The best approximate value function $\tilde{V}$ is that function in $\mathcal{F}$ (or equivalently $\mathcal{F}^{T}$) that minimizes the above expected norm difference.
Suppose we are given just $N$ samples $\bar{s}_{1},\dots,\bar{s}_{N}$ from a Markov chain induced by policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$, and corresponding rewards $r_{1},\dots,r_{N}$. The next best thing is to minimize the empirical norm difference at these points, defined as
$\left|\left|f\right|\right|_{N}=\left(\frac{1}{N}\sum_{t=1}^{N}f(x_{t})^{2}\right)^{\frac{1}{2}}.$
The above function defines a norm, along with a corresponding inner product, in a new inner product space $\mathcal{F}_{N}$, the space of all ($N$) values of the functions at the given sample points. This can be considered as a subset of $\mathbb{R}^{N}$ with the obvious inner product.
$\mathcal{F}_{N}$ =$=\left\{\left(f(\bar{s}_{1}),\dots,f(\bar{s}_{N})\right):f\in\mathcal{F}\right\}$
=$=\left\{\left(f(s^{T}_{1}),\dots,f(s^{T}_{N})\right):f\in\mathcal{F}^{T}\right\}$
=$=\left\{\Phi\alpha:\alpha\in\mathbb{R}^{d}\right\},$
where $\Phi=\left(\phi(\bar{s}_{1}^{T}),\dots,\phi(\bar{s}_{N}^{T})\right)_{N\times d}$, and $\bar{s}_{i}^{T}=f^{T}_{trunc}(\bar{s})$.
The true value function $V$ is a fixed point of the Bellman operator $\mathcal{T}$ in $\bar{S}$. From the given $N$ points, we can define our approximate value function as the fixed point of the projected pathwise Bellman operator $\hat{\Pi}\hat{\mathcal{T}}$, where $\hat{T}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the pathwise Bellman operator, with $\hat{T}(x)$ being defined for $x\in\mathbb{R}^{N}$ as
$\left(\hat{T}(x)\right)_{i}=\begin{cases}r_{i}+\gamma x_{i+1}&\text{ for }i<N\\ r_{i}&\text{ for }i=N\end{cases},$
and $\hat{\Pi}:\mathbb{R}^{N}\rightarrow\mathcal{F}_{N}$ is the projection operator onto the valid function approximated values on $x_{1:N}$, i.e,
$\hat{\Pi}(x)=\mathop{\mathrm{argmin}}\limits_{z\in\mathcal{F}_{N}}\left|\left|x-z\right|\right|_{N}^{2}.$
This also be written as $\hat{T}=r+\gamma\hat{P}$, where $r=(r_{t})_{t\in[N]}$ and $\hat{P}$ is the path-wise transition operator. $\hat{T}$ is a $\gamma$-contraction in ||.$||.||_{N}$ as well as ||.$||.||_{2}$, and the projection operator $\hat{\Pi}$ onto $\mathcal{F}_{N}$ is non-expansive since $\mathcal{F}$ is convex.
This means that $\hat{\Pi}\hat{T}$ is a $\gamma$-contraction on $\mathbb{R}^{N}$ and has a unique fixed point $\hat{v}\in\mathcal{F}_{N}$, which we define as our approximation of the value function values at these points. Corresponding to this point, there exists function $\hat{V}\in\mathcal{F}$ such that $\hat{v}=\left(\hat{V}(\hat{s}_{t})\right)_{t\in[N]}$.
Given samples $\bar{s}_{1},\dots,\bar{s_{N}}$, the approximate value function is defined as
$\hat{V}=\sum_{i\in[d]}\hat{\alpha}_{i}\varphi_{i},\text{ where }\hat{\alpha}=\left[\Phi^{T}\left(I-\gamma\hat{P}\right)\right]^{+}\Phi^{T}r.$
This $\hat{V}$ is considered the output of the approximate policy evaluation step of the policy iteration algorithm. For determining the overall error of this estimate, the intermediate task is to bound the difference between this and the true solution on these data points, i.e, $||V-\hat{V}||_{N}$, as a function of the number of data points and the complexity of the function class under consideration.
Let $v=\left(V(\bar{s}_{t})\right)_{t\in[N]}$ just as $\hat{v}=\left(\hat{V}(\bar{s}_{t})\right)_{t\in[N]}$. Following Lazaric et al. (2012),
$||V-\hat{V}||_{N}=||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}.$
The first term is the approximation error due to having a restricted function space and is unavoidable. The second term also includes the estimation error due to using the pathwise Bellman operator instead of the actual Bellman operator, and is a function of the number of samples used. It can be written as $||\hat{\Pi}\xi||_{N}$, where $\xi=\hat{\mathcal{T}}v-v$ is the estimation error. Now, for $t<N$,
$\xi_{t}$ $=r_{t}+\gamma V(\bar{s}_{t+1})-V(\bar{s}_{t})$
$=r(s_{t},\pi(\bar{s}_{t}),s_{t+1})+\gamma V(\bar{s}_{t+1})-\mathsf{E}_{\bar{S}_{t+1}}\left[r(s_{t},\pi(\bar{s}_{t},\bar{S}_{t+1})+\gamma V(\bar{S}_{t+1})\right]$
$=r(s_{t},\pi(\bar{s}_{t}),s_{t+1})-\mathsf{E}_{S_{t+1}\sim Q_{x_{:t}}(.|s_{t},\pi(\bar{s}_{t}))}\left[r(s_{t},\pi(\bar{s}_{t}),S_{t+1})\right]$
$\qquad+\gamma\left[V(\bar{s}_{t+1})-\mathsf{E}_{\bar{S}_{t+1}\sim Q,Q^{X}_{x_{:t}}(.|s_{t},\pi(\bar{s}_{t}))}V(\bar{S}_{t+1})\right],$
and for $t=N$,
$\xi_{t}$ $=r(s_{t},\pi(\bar{s}_{t}),s_{t+1})-\mathsf{E}_{S_{t+1}\sim Q_{x_{:t}}(.|s_{t},\pi(\bar{s}_{t}))}\left[r(s_{t},\pi(\bar{s}_{t}),S_{t+1})\right]$
$\qquad-\gamma\mathsf{E}_{\bar{S}_{t+1}\sim Q,Q^{X}_{x_{:t}}(.|s_{t},\pi(\bar{s}_{t}))}\left[V(\bar{S}_{t+1})\right].$
$\xi_{t}$ are functions of $\bar{s}_{t}$, that are samples from the Markov chain induced by the policy $\pi$ in the augmented MDP $\bar{M}$. Considering them as a random variables, $\left(\xi_{t}\varphi_{t}(\bar{s_{t}})\right)_{1\leq t<N}$ is a Martingale difference sequence w.r.t $\bar{s}_{t}$, with each element being zero mean and bounded by $\frac{L}{1-\gamma}$. So, applying Azuma’s inequality to $(\xi_{t}\varphi_{i}(\bar{s})_{t})_{1\leq t<N}$ gives, for each $i\in[d]$,
$\mathsf{P}\left(\left|\sum_{t=1}^{N-1}\xi_{t}\varphi_{i}(s_{t})\right|\geq\epsilon\right)<2\exp\left(-\frac{\epsilon^{2}(1-\gamma)^{2}}{2(N-1)L^{2}}\right).$
Letting the right hand side be $\frac{\delta}{d}$, with probability at least $1-\delta$,
$\left|\sum_{t=1}^{N-1}\xi_{t}\varphi_{i}(s_{t})\right|<\frac{L}{1-\gamma}\sqrt{2(N-1)\log(2d/\delta)}\quad\forall i\in[d].$
Since $|\xi_{n}\varphi_{i}(\bar{s}_{t})|<\frac{L}{1-\gamma}$ always, we have,
$\left|\sum_{t=1}^{N-1}\xi_{t}\varphi_{i}(s_{t})\right|<\frac{L}{1-\gamma}\left(\sqrt{2(N-1)\log(2d/\delta)}+1\right)\;\forall i\in[d]\text{ w.p. }\geq 1-\delta.$
Following Lazaric et al. (2012), this can be used to bound $||\hat{\Pi}\xi_{n}||_{N}$ as
$\left|\left|\hat{\Pi}\hat{\mathcal{T}}v-\hat{\Pi}v\right|\right|_{N}\leq\frac{1}{N}\sqrt{\frac{d}{\nu_{n}}}\max_{i}\left|\sum_{t=1}^{N}\xi_{t}\varphi_{i}(s_{t})\right|,$
where $\nu_{n}$ is the smallest positive eigenvalue of the Gram matrix $\Phi^{T}\Phi$.
Substituting these terms in the original inequality, we obtain
$||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}$
$\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu_{N}}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right).$ (2)
The above identity is almost identical to the one in Lazaric et al. (2012), the reason being that we are still operating in the range space of the basis functions restricted to a finite set of states and so the structure of the MDP in our setting does not affect the analysis. Since the proof follows a similar path, we have only included those steps where there is a discrepancy, which are mainly due to the noise terms taken in the martingale difference sequence. In our setting, these noise terms also include stochasticity due to the reward in addition to the transition function, leading to a difference in the final expression.
While the final inequality is quite similar, we shall see that when considering the expected error in the augmented state space, we can bring out the properties of our setting by splitting the expected error in terms of error due to function truncation and the inherent error due to linear function approximation.
7.2. Expected error
The above inequality bounds the average error of the estimated value function evaluated only at those points obtained as a sample from the environment. The expected error everywhere with respect to some distribution can be bounded in terms of the above bound considering assumptions on the function class and convergence rate of the Markov chain induced by the policy under consideration.
For linear function spaces, and assuming that the policy induces a $\beta$-mixing Markov chain, following the proof of Lazaric et al. (2012) gives the following generalization bound for policy evaluation. We state the theorem in a simple form here for clarity. The full details of the theorem, as well as supporting Lemmas, are given in the Appendix.
Theorem 7.1.
Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using policy $\pi$ in the MDP $\mathcal{M}_{X}$. Suppose sufficiently large $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$. Let $\nu$ be a lower bound on the eigenvalues of the sample-based Gram matrix that hold with probability $41-\delta/$ 4. Then, with probability at least $1-\delta$,
$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(8d/\delta)}{N}}+\frac{1}{N}\right)+\bar{\epsilon}(N,d,\delta).$
Proof.
Following proof of Theorem $5$ of Lazaric et al. (2012),
$2||\hat{V}-V||_{N}\geq 2||\tilde{V}-V||_{N}\geq||\tilde{V}-V||_{\rho}-\epsilon_{1},$
where the first inequality is because of truncation, and the second inequality is a generalization Lemma for Markov chains, and so
$||\tilde{V}-V||_{\rho}$ $\leq 2||\hat{V}-V||_{N}+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\hat{\Pi}V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{\rho}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-\Pi^{trunc}V||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}||_{\rho}+||V^{\epsilon}-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}.$
The second inequality is due to Equation 2, the third is by the definition of $\hat{\Pi}$, and the fourth is the generalization bound in the opposite direction as the first. The fifth is just the triangle inequality.
The sixth inequality is by definition of the $\Pi^{trunc}$ operator, where $V^{\epsilon}$ is the ‘truncated’ value function of an $\epsilon$-suboptimal policy that depends only on the past $T$ events. This is guaranteed from the proof of Theorem 5.1 for
$\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
The seventh inequality is due to the triangle inequality by adding and subtracting the actual value function $V^{\epsilon}$ of the $\epsilon$-suboptimal policy. The final eighth inequality is due to Lemma 6.1 and the value of $\epsilon$.
Both these generalization bounds happen with probability at least $1-\delta^{\prime}$ each. Further, with the lower bound $\nu$ on $\nu_{N}$ holding with probability at least $1-\delta^{\prime}$, the final bound holds with probability at least $1-\delta=1-4\delta^{\prime}$. ∎
The generalization terms due to linear function approximation using Markov samples remain essentially the same as in Lazaric et al. (2012). The main difference is the first extra term in the error that is induced due to truncating the augmented state space by discarding the features of old events. The amount of error introduced due to this approximation depends on the amount of influence exerted by old events on the current events as well as state transition kernel.
Remarks on the proof
The Lemmas used in the proof can be taken from Lazaric et al. (2012) without any modifications because, even if some of the results, specifically the results on the bound on the covering numbers for linear function spaces taken from Györfi et al. (2002), require the domain of the basis functions to be finite dimensional, the analysis takes place in the co-domain and so can be directly applied to our setting. We state the complete theorem along with the Lemmas in Appendix B.
8. Experiments
We conducted experiments for solving a non-stationary variant of classic Pendulum-v1 environment in the control suite of OpenAI Gym. The task consists of applying torque to a pendulum to swing it and keep it upright. We considered a fixed neural-network policy that has been pre-trained partially using DDPG. We then used pathwise LSTD with linear function approximation for evaluating this policy as described in Section 7. The features considered are the standard cosine and sine of the angle and the angular velocity, with additional features being the angle itself along with the squares of all these features, making the state $8$ dimensional.
Figure 2. Plots showing the dependence of the Mean Squared Error of the value function learned using pathwise LSTD as a function of the rate of decay of the exogenous Hawkes process, the horizon of past events considered and the number of samples used.
For the non-stationarity, we consider the discrete-time Gaussian marked Hawkes process described in Section 4.2. The intensity due to an event is decayed as $\alpha=e^{-\lambda_{\alpha}t}$ with $\lambda_{\alpha}=1.0$, and the effect on the event mark decays as $\beta_{t}=\frac{1}{1+t^{2}}$. The events are added to the torque applied to the pendulum. Figure 2 shows the expected error of the learned value function as a function of the number of samples, the event horizon $T$, and the rate $\lambda_{\alpha}$ of decay of the Hawkes process.
For a fixed event horizon $5$, the error reduces as the number of samples increases, which is quite intuitive. For a fixed number of samples $10000$, the average error initially reduces as the event horizon increases, corresponding to the first term in the sample complexity of the total variation induced by events older than the event horizon. After a certain stage, increasing the event horizon results in a gradual increase in the error. This is because the events older than the horizon now contribute a negligible influence on the current dynamics, while the dimensionality of the features is increasing due to the inclusion of more past events in the features, resulting in a slightly greater expected error. The third plot shows the dependence on $\lambda_{\alpha}$, the rate of decay of $\alpha_{t}$. Faster decay results in lesser non-stationarity and lower approximation error due to the first term in the sample complexity upper bound. Due to high variation in the results obtained, we conducted $20$ trials and plotted the median of the Mean Squared Errors for the experiments w.r.t decay rate and number of samples. The error bars indicate the range of $40$-$60$ percentile versus the decay rate and $20$-$80$ percentile for the other two.
9. Discussion and Future Directions
State space
The theorem proposed is general and relies on learning functions in an augmented state space that blows up in the presence of long-term dependencies due to the external events. For a practical algorithm, it is necessary to compress this augmented state space to something more manageable.
This can be done implicitly by using function approximators such as recurrent neural networks or LSTMs, or explicitly by learning compressed state representations using mutual information bottlenecks etc. While these are practical methods, it is also an open problem on how to theoretically analyze such a learnt compressed state representation.
Generalization
The sample complexity analysis in Section 7 hinges on generalizing sample error of policy evaluation to an expected error over the entire state space. Such generalization generally requires assumptions on the distribution over states.
In the theorem stated by us, the assumption was on the rate of convergence of the Markov chain induced from the augmented MDP by the policy under evaluation, which is common in the reinforcement learning literature. However, in our case, the specific structure of the state space, i.e, separation of the augmented state space into states and events with different characteristics, may lead itself to a different types of assumptions. For example, the recurrence and aperiodicity of the Markov chain may not be a good assumption, but a stationarity assumption on just the event sequence might be justified.
Due to this, the generalizability of the value function in our setting can be analyzed in a more specific way that takes into account this exogenous structure.
Policy Improvement
The policy improvement step involves optimizing over the action space based on the expected one-step reward and the value function at the next time step, which requires knowledge of the model. Thus, in the case of an unknown transition and reward model, it is important to investigate the sample complexity of learning the models from samples.
Appendix A Proofs of lemmas
A.1. Proof of Lemma  5.2
We adapt the proof of Theorem 1 from Xiao et al. (2019).
We wish to study the value function of the policy in two different MDPs $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$. For convenience, without any loss in generality, we can extend the state space of $\mathcal{M}^{(T)}$ to that of $\mathcal{M}_{X}$ by including the information of every past event in the state, without changing the transition function.
We define another class of auxiliary MDPs $\mathcal{M}^{(T)}_{h}$, $h\in 0\cup\mathbb{N}$, in which when starting from initial (augmented) state $\bar{s}\in\bar{S}$, the transitions follow the transition kernel of $\mathcal{M}_{X}$ for $h$ time steps and then the transition kernel of $\mathcal{M}^{(T)}$ for all transitions thereafter. The value functions induced by $\pi$ in $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$ can be related using these new intermediate MDPs, which interpolate between those two environments.
Since only one policy is under discussion, the $\pi$ in $V^{\pi}$ is omitted for the rest of this proof for simplicity of notation. $V$ denotes the value function of the policy in $\mathcal{M}_{X}$, $V^{(T)}$ denotes its value in $\mathcal{M}^{(T)}$, and $V_{h}^{(T)}$ in $\mathcal{M}^{(T)}_{h}$.
The following relations follow from the definitions of the auxiliary and intermediate MDPs and the boundedness of $r$:
$V^{(T)}=V_{0}^{(T)},\quad V=\lim_{h\rightarrow\infty}V^{(T)}_{h}$
and so,
$V-V^{(T)}$ $=\lim_{H\rightarrow\infty}V_{H}^{(T)}-V_{0}$
$=\lim_{H\rightarrow\infty}\left(V_{H}^{(T)}-V_{0}\right)$
$=\lim_{H\rightarrow\infty}\sum_{h=0}^{H-1}\left(V_{h+1}^{(T)}-V_{h}^{(T)}\right)$
$=\sum_{h=0}^{\infty}\left(V_{h+1}^{(T)}-V_{h}^{(T)}\right).$
That is, the entire difference can be characterized in terms of differences in $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ for each $h$. Now for any $s\in\bar{S}$,
$V_{h}^{(T)}(s)$ $=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]$
$\quad+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\quad+\sum_{t=h+1}^{\infty}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t-h}^{(T)}\circ p_{h}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p^{(T)}(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right],$
where $p_{t}(.|s)$ denotes a transition of $t$ steps in the environment $\mathcal{M}_{X}$ starting from $s$, a superscript (T) denotes the corresponding object in $\mathcal{M}^{(T)}$, and the policy $\pi$ is implicit whenever it is not explicitly mentioned in any expression.
This can also be written as
$V_{h}^{(T)}(s)=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h}).$
Using these two representations for $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ respectively gives
$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{h+1}\sim p(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p_{h+1}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
$-\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p^{(T)}\circ p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
$=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})\bigg{]}$
$+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h},a_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}V^{(T)}(s_{h+1})$
$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}V^{(T)}(s_{h+1})\bigg{]}.$
In the above equations, $s,s_{t}$ etc., denote the augmented state, which consists of the state and event mark history till then. Since the current state and the current event mark are independent and depend only on the previous events, the transition kernel can be factored as two independent distributions over $S$ and $\mathcal{X}$. So, replacing the augmented state $s\in\bar{S}$ with the explicit representation $(s,H)$, where $s\in S$ and $H\in\mathcal{X}^{\infty}$ gives, for any bounded measurable function $f$ on $\mathbb{K}\times\mathcal{X}^{\infty}$,
$\mathsf{E}_{(s_{t+1},H_{t+1})\sim p(.|s_{t},H_{t},a_{t})}f(s_{t+1},H_{t+1})-\mathsf{E}_{(s_{t+1},H_{t+1})\sim p^{(T)}(.|s_{t},H_{t},a_{t})}f(s_{t+1},H_{t+1})$
$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$+\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
$\leq\mathsf{E}_{\begin{subarray}{c}X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}\left[\left|\left|f(.,X_{t+1},H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q_{H_{t}}(.|s_{t},a_{t}),Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$+\mathsf{E}_{s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})}\left[\left|\left|f(s_{t+1},.,H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q^{X}_{H_{t}}(.|s_{t},a_{t}),Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
$\leq||f||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+\sum_{t=T+1}^{\infty}N_{t}\right).$
The second to last inequality is a result of Total Variation Distance being an integral probability metric (Sriperumbudur et al., 2009). Using this expression in the previous equation gives
$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $\leq\gamma^{h}||r||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\qquad\qquad+\gamma^{h+1}||V^{(T)}||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\leq\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
This gives the final result,
$||V-V^{(T)}||_{\infty}$ $\leq\sum_{h=1}^{\infty}\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$=\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
A.2. Proof of Lemma 6.1
$V^{\pi}((s,x_{0:\infty}))$ $=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:\infty}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left[r(s,a,s^{\prime})+V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))\right],$
$V^{\pi}((s,x_{0:T},0))$ $=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left[r(s,a,s^{\prime})+V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right].$
Taking the difference between these two expressions, and adding and subtracting the additional term
$\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))$
gives
$\left|V^{\pi}((s,x_{0:\infty}))-V^{\pi}((s,x_{0:T},0))\right|$
$\leq||r||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}\right)+||V^{\pi}||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+N_{t}\right)$
$\qquad\qquad+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left|V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))-V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right|.$
In the above inequality, the left-hand side is the difference between two value functions whose first $T$ events are the same and differ everywhere else, but the right-hand side has a discounted term in which the first $T+1$ terms coincide. So, repeating the same procedure results in value function differences at states that increasingly coincide.
Since each repetition adds a $\gamma$ factor to the term, which itself is bounded by $\frac{1}{1-\gamma}$ because of bounded rewards, the resulting series converges, giving us the following bound:
$\sup_{\bar{s}=(s,x_{0:\infty})}\left|V^{\pi}((s,x_{0:\infty}))-V^{\pi}((s,x_{0:T},0))\right|$
$\leq\sum_{t=T+1}^{\infty}\gamma^{t-T-1}\left[||r||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}\right)+||V||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}+N_{t^{\prime}}\right)\right]$
$\leq\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}\sum_{t^{\prime}=t}^{\infty}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$\leq\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\sum_{t=0}^{T-t^{\prime}-1}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
$=\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\frac{1-\gamma^{t^{\prime}-T}}{1-\gamma}(M_{t^{\prime}}+N_{t^{\prime}})\leq\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
Appendix B Details of Theorem 7.1
Intuitively, bounding the expected error of the estimated value function based on the error at just a few samples requires the samples to be close to the stationary distribution, and the function values at the samples to be close to their expected values.
The first requirement is formalized in terms of the rate of mixing of the Markov chain induced from the MDP by the policy.
Definition 1.
A Markov chain $\mathcal{M}=\left(X_{t}\right)_{t\geq 1}$ is said to be exponentially $\beta$-mixing with parameters $\bar{\beta},b,\kappa$ if its mixing coefficients
$\beta_{i}=\sup_{t\geq 1}\left[\sup_{B\in\sigma\left(X_{t+i},\dots\right)}\left|\mathsf{P}\left(B|X_{1},\dots,X_{t}\right)-\mathsf{P}(B)\right|\right]$
satisfy $\beta_{i}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right)$.
When such an exponentially mixing Markov chain is also ergodic and and aperiodic with stationary distribution $\rho$, for any initial distribution $\lambda$, there is a bound on the following total variation: (Definition 21 of Lazaric et al. (2012))
$\left|\left|\int_{\mathcal{X}}\lambda(dx)\mathsf{P}\left(.|x\right)-\rho(.)\right|\right|_{\mathsf{TV}}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right).$
This helps in bounding the difference between the norm of a function in the stationary distribution and the empirical distribution defined using a few samples.
Lemma B.1 (Generalization Lemma for Markov chains, Lemma 24 of Lazaric et al. (2012)).
Let $\tilde{\mathcal{F}}$ be the $d$-dimensional class of linear functions truncated at threshold $B$. Let $\left(X_{t}\right)_{t\in[n]}$ be a sequence of samples from an ergodic, aperiodic exponentially $\beta$-mixing Markov chain with parameters $\bar{\beta},b,\kappa$, arbitrary initial distribution, and stationary distribution $\rho$. If the first $\tilde{n}$ samples are discarded and only the last $n-\tilde{n}$ samples are used, then with probability at least $1-\delta$, the empirical and $l_{2}(\rho)$ norm of the any function $\tilde{f}\in\tilde{\mathcal{F}}$ are related as,
$||\tilde{f}||-2||\tilde{f}||_{X_{1:n}}$ $\leq\epsilon(\delta),$
$||\tilde{f}||_{X_{1:n}}-2\sqrt{2}||\tilde{f}||$ $\leq\epsilon(\delta),$
$\text{for }\epsilon(\delta)=12B\sqrt{\frac{2\Lambda(n-\tilde{n},d,\delta)}{n-\tilde{n}}\max\left\{\frac{\Lambda(n-\tilde{n},d,\delta)}{1},1\right\}^{\frac{1}{\kappa}}},$
$\Lambda(n,d,\delta)=2(d+1)\log n+\log\frac{\epsilon}{\delta}+\log^{+}\left(\max\left\{16(6e)^{2(d+1)},\bar{\beta}\right\}\right),$
[…]
The above lemma is essentially a generalization bound for truncated linear functions operating on samples from an exponentially mixing Markov chain, and is used in the proof of Theorem 7.1 to translate the bound on the empirical error to a bound on the expected error in the stationary distribution of the Markov chain induced by the policy under evaluation.
The above bounds hold for every member of the $d$-dimensional class of functions, hence the dependence of $\epsilon$ on $d$ through $\Lambda(n,d,\delta)$. In contrast, when there is just one truncated linear function under consideration, the above bound holds, but with
$\Lambda(n,\delta)=\log\frac{\epsilon}{\delta}+\log\left(\max\left\{6,n\bar{\beta}\right\}\right).$
Another issue for generalizing the sample-based bound in Equation 2 to the entire state space is its dependence on the eigenvalues of the sample Gram matrix. For tackling this problem, Lazaric et al. (2012) derived the following probabilistic lower bound on the smallest eigenvalue of the sample-based Gram matrix as a function of the smallest eigenvalue of the Gram matrix $G=\mathsf{E}_{x\sim\rho}\left[\phi(x)\phi(x)^{T}\right]$.
Lemma B.2 (Lemma 4 of Lazaric et al. (2012)).
Let $\omega>0$ be the smallest eigenvalue of the Gram matrix defined above. If the data generating process is an exponentially $\beta$-mixing Markov chain with parameters $\bar{\beta},b,\kappa$, then the smallest eigenvalue $\nu_{n}$ of the sample-based Gram matrix constructed using $n$ samples satisfies
$\sqrt{\nu_{n}}\geq\sqrt{\nu}=\frac{\sqrt{\omega}}{2}-6L\sqrt{\frac{2\Lambda(n,d,\delta)}{n}\max\left\{\frac{\Lambda(n,d,\delta)}{b},1\right\}^{\frac{1}{\kappa}}}>0,$
provided the number of samples $n$ satisfies
$n>\frac{288L^{2}\Lambda(n,d,\delta)}{\omega}\max\left\{\frac{\Lambda(n,d,\delta)}{b},1\right\}^{\frac{1}{\kappa}},$
where
$\Lambda(n,d,\delta)=2(d+1)\log n+\log\frac{e}{\delta}+\log^{+}\left(\max\left\{18(6e)^{2(d+1)},\bar{\beta}\right\}\right)$
comes from a simpler version of Lemma B.1 without the additional initial $\tilde{n}$ samples for burn-in of the Markov chain.
The generalization Lemma B.1, combined with the lower bound on the eigenvalues and our results in Section 6, give rise to the proof of the following result on the sample complexity of policy evaluation, detailed in Section 7.
B.1. Theorem 7.1
Assume that the policy $\pi$ induces from the MDP $\mathcal{M}_{X}$ a $\beta$-mixing Markov chain with parameters $\beta$, $b$, $\kappa$. Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using the policy $\pi$. Suppose the first $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$. Then, provided the number of discarded samples is $\tilde{N}=\left(\frac{1}{b}\log\frac{2e\bar{\beta}n}{\delta}\right)^{\frac{1}{\kappa}}$, with probability at least $1-\delta$,
$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right)$
$\qquad\qquad+\epsilon_{1}(N,d,\delta)+2\sqrt{2}\epsilon_{2}(N,\delta),$
where $\text{where }\epsilon_{1}(N,d,\delta)$ $=\frac{24}{1-\gamma}\sqrt{\frac{2\Lambda_{1}(N,d,\delta/4)}{N}\max\left\{\frac{\Lambda_{1}(N,d,\delta/4)}{b},1\right\}^{\frac{1}{\kappa}}},$
$\epsilon_{2}(N,\delta)$ $=12\left(\frac{1}{1-\gamma}+L||\alpha^{*}||\right)\sqrt{\frac{2\Lambda_{2}(N,\delta/4)}{N}\max\left\{\frac{\Lambda_{2}(N,\delta/4)}{b},1\right\}^{\frac{1}{\kappa}}},$
with $\Lambda_{1}(N,d,\delta)$ and $\Lambda_{2}(N,\delta)$ as defined in Lemma B.1, and $\alpha^{*}$ such that $\Pi V=\sum_{i}\alpha^{*}_{i}\varphi_{i}$.

### v3

@@
 To address this problem, we study Markov Decision Processes (MDP) under the influence of an external temporal process.
-We formalize this notion and discuss conditions under which the problem becomes tractable with suitable solutions.
+First, we formalize this notion and derive conditions under which the problem becomes tractable with suitable solutions.
 We propose a policy iteration algorithm to solve this problem and theoretically analyze its performance.
-We derive results on the sample complexity of the algorithm and study its dependency on the extent of non-stationarity of the environment.
-We then conduct experiments to illustrate our results in a classic control environment.
+Our analysis addresses the non-stationarity present in the MDP as a result of non-Markovian events, necessitating the formulation of policies that are contingent upon both the current state and a history of prior events.
+Additionally, we derive insights regarding the sample complexity of the algorithm and incorporate factors that define the exogenous temporal process into the established bounds.
+Finally, we perform experiments to demonstrate our findings within a traditional control environment.
 1. Introduction
-Reinforcement Learning (RL) provides a framework for sequential decision-making problems, wherein an agent has to make decisions at each time step so that a given objective is achieved.
-A huge amount of research has been done in this field in recent years, resulting in remarkable progress in various applications like robotics, finance, and even video games (Zhao et al., 2020; Charpentier et al., 2021; Mnih et al., 2015).
-Central to RL is the mathematical framework of Markov Decision Processes (MDPs).
-In an MDP, an agent is assumed to be in an ‘environment’, and at each time step, the agent has to take a decision or an ‘action’.
-The state of the environment changes stochastically in a Markov manner.
-It depends only on the current state and the agent’s current action.
-While most RL research models the agent’s environment as an MDP, this paper argues that this is a very simplistic model of the real world, as one assumes that the state of a Markov Decision Process might constantly be changing, but the rules of the state transition are still fixed.
+Reinforcement Learning (RL) serves as a foundational framework for addressing sequential decision-making problems.
+In recent years, extensive research in this domain has led to significant advancements across various fields, including robotics, finance, and navigation.
+At the core of RL lies the mathematical framework of Markov Decision Processes (MDPs).
+In this framework, an agent interacts with an ‘environment’, where, at each time step, the agent is required to make a decision to take an ‘action.’ The state of the environment changes stochastically in accordance with the Markov property, relying solely on the present state and the action taken by the agent.
+While solving the various sequential decision-making problems, in most RL research, it is assumed that while the state of an MDP may be constantly changing, the rules governing state transitions remain unchanged.
 However, in practical applications, external events may influence the agent’s environment and change its dynamics.
-For example, consider the problem of portfolio management in finance.
-An agent has access to a state consisting of various fundamental and technical indicators of a select number of financial instruments and has to make decisions, sequentially, on whether to hold on to an instrument, sell it, or short it, etc.
-But the way the prices of these instruments move may be affected by external forces, such as government decisions like rate hikes or sudden market movements caused by events occurring in a completely different sector of the economy not being considered by the agent.
+For example, in the context of portfolio management in finance, an agent has access to a state consisting of various fundamental and technical indicators of a select number of financial instruments and has to make decisions sequentially on whether to hold on to an instrument, sell it, short it, etc.
+However, the price movements of these instruments may be affected by external forces, such as government decisions like rate hikes or sudden market movements caused by events occurring in a completely different sector of the economy that are not being considered by the agent.
 Such an external event might exert lasting influence over different time scales depending on the kind of event that occurred.
 A rate cut might affect stock prices for a few months, while a flash crash might cease its effect in a few hours.
+An example illustrating the persistent non-stationarity effect of perturbations due to exogenous events is depicted in Figure 1.
 As another example, consider a navigation problem in which an agent has control of a robot and needs to travel from one point to another.
 While the agent can learn to do this task in ideal conditions, the optimal actions taken by the agent might depend on external conditions such as rain, sudden traffic changes, or human movement patterns.
-So, while an agent can learn to achieve goals in an ideal environment, it cannot be oblivious to external changes that influence the environment dynamics differently.
-To tackle this problem, we consider a Markov Decision Process that is perturbed by an external discrete-time temporal process.
-This external process affects the dynamics of the MDP in a non-Markovian fashion, resulting in a non-stationary problem.
+Hence, while an agent can learn to achieve goals in an ideal environment, it cannot be oblivious to external changes that influence the environment dynamics differently.
 Figure 1.
-An illustration of the lasting non-stationary effect of perturbations due to exogenous events.
 The plot shows the prices of two assets following a Geometric Brownian motion.
 A few exogenous Gaussian shocks to one asset forces it to diverge significantly from the other over the long term, with a markedly different pattern of evolution long after the effect of the exogenous event decays.
+In the literature, non-stationarity settings in RL have been studied, albeit considering only very special cases, mostly providing practical solutions.
+Some model non-stationarity as piece-wise stationarity (Padakandla et al., 2020; Alegre et al., 2021; Hadoux et al., 2014), in which the environment remains unchanged except during a select few time steps called change points, during which it can change arbitrarily.
+Few works consider drifting environments wherein the transition and reward functions can vary at any time step while satisfying some constraints to make the problem tractable (Lecarpentier and Rachelson, 2019; Cheung et al., 2020).
+Hallak et al.
+(2015) study Contextual MDPs, in which the environment dynamics change based on externally specified “contexts” that are episodic and possibly unknown chosen from among a finite set of known values.
+Tennenholtz et al.
+(2023) extend this setting to handle dynamic contexts that are known to the agent, change at each time step, and influence the state transition distribution.
+However, their contexts are not exogenous and come from a finite set of possible vectors.
+In contrast to the above existing literature, we study this problem in a general setting, where a Markov Decision Process is perturbed by an external discrete-time temporal process.
+This external process affects the dynamics of the MDP in a non-Markovian fashion, resulting in a non-stationary problem.
+This process, being exogenous to the decision process, does not depend on the state of the environment under consideration or the actions of the agent.
+In this paper, we study conditions that the exogenous process has to satisfy, along with its effect on the MDP, such that almost optimal solutions exist that are tractable in some sense.
+We formulate the variants of standard reinforcement learning algorithms that consider above-mentioned non-stationarity problem and theoretically analyze the properties of these algorithms.
 Contributions
   1. (1)
-    We formulate a non-stationary reinforcement learning problem consisting of an MDP perturbed by an external discrete-time temporal process that affects the dynamics of the MDP in a non-Markovian fashion.
-We describe sufficient conditions under which such a problem has a well-defined solution.
+    We present a non-stationary reinforcement learning formulation for an MDP under the influence of an external discrete-time temporal process, which alters the MDP dynamics in a non-Markovian manner.
+We outline the necessary conditions that ensure the existence of a well-defined solution for this problem.
   2. (2)
-    We derive conditions under which a solution can be determined approximately by an agent based on just the current state and a finite history of past events.
-This is possible when the perturbations caused by events older than $t$ time steps on the MDP transition dynamics and the event process itself are bounded in total variation by $M_{t}$ and $N_{t}$ respectively, and $\sum_{t}M_{t},\sum_{t}N_{t}$ are convergent series.
+    We establish criteria under which an approximate solution can be determined using only the current state and finite history of past events.
+We show that this is possible when the perturbations caused by events older than $t$ time steps on the MDP transition dynamics and the event process itself are bounded in total variation by $M_{t}$ and $N_{t}$ respectively, and $\sum_{t}M_{t},\sum_{t}N_{t}$ are convergent series.
 We analyze the trade-off between the performance of the agent and the length of history considered.
   3. (3)
-    We propose a policy iteration algorithm to solve this problem and theoretically analyze its behavior.
-We show that policy improvement occurs in regions of the state space where the Bellman error is not too low based on $M_{t}$ and $N_{t}$.
-Higher the approximation error due to non-stationarity, smaller the region where the policy is guaranteed to improve.
+    We propose a policy iteration algorithm and theoretically analyze its behavior.
+We show that policy improvement occurs in regions of the state space where the Bellman error is “not too low” based on $M_{t}$ and $N_{t}$.
+Based on our analysis, we observe that the higher the approximation error due to non-stationarity, the smaller the region where the policy is guaranteed to improve.
   4. (4)
@@
     We further explain our theoretical results using the discrete-time Hawkes process and conduct experiments in a non-stationary pendulum environment to validate our results.
-2. Related Work
-While most Reinforcement Learning research considers a stationary Markov Decision Process, there has been some work done on tackling various kinds of non-stationarities that might affect the MDP.
-Many works consider environments that are piece-wise stationary, i.e., the transition and reward functions remain unchanged except during a select few time steps called change points, during which they can change arbitrarily.
-Different works consider different assumptions on prior knowledge about the nature of the change points etc., available to the agent.
-Padakandla et al.
-(2020) assume that the transition and reward functions belong to a fixed set of ‘contexts’ between change points.
-They propose a model-free algorithm that uses a change point detection algorithm that determines when a change occurs by analyzing the state-reward-next state tuples obtained from the environment.
-However, their algorithm assumes knowledge of the pattern of changes in the context.
-Alegre et al.
-(2021) propose a model-based algorithm that can deal with arbitrarily different types of contexts and continuous state and action spaces.
-Other works consider drifting environments wherein the transition and reward functions can vary at any time step, while satisfying some constraints to make the problem tractable.
-Lecarpentier and Rachelson (2019) consider a finite horizon MDP in which the transition dynamics and reward function evolve such that they are Lipschitz continuous with respect to time.
-Cheung et al.
-(2020) consider such a setting in a finite MDP where the amount of total change in the environment, referred to as the variation budget, is known to the agent.
-They propose an optimistic Value Iteration-based algorithm and prove guarantees on the upper bound of its dynamic regret.
-Wei and Luo (2021) propose a black-box reduction that turns any suitable algorithm with optimal performance in a near-stationary environment into an algorithm that achieves optimal dynamic regret in a non-stationary environment.
-Hadoux et al.
-(2014) formalized a class of piecewise stationary MDPs called Hidden-Semi-Markov-Model MDPs where the transition from one hidden stationary MDP mode to another is semi-Markov in nature.
-To solve this problem, they adapt the Partially Observable Monte Carlo Planning (POMCP) algorithm used for solving Partially Observable Markov Decision Processes (POMDPs).
-Some algorithms assume additional structure on the nature of non-stationarity.
-Feng et al.
-(2022) propose Factored Non-stationary MDPs (FN-MDPs), where the transition dynamics are defined in terms of a Dynamic Bayesian Network over the states, actions, rewards, and latent change factors that cause non-stationarity.
-Here, both the states and the latent change factors have transition dynamics obeying this causal graph.
-They propose using a Variational Auto-Encoder (VAE) framework to learn transition dynamics and present results in MuJoCo and robotic manipulation tasks.
-Another factorization setting with a resemblance to ours has been proposed by Efroni et al.
-(2022) called the Exogenous MDP, wherein the environment state is divided into two parts, an endogenous state which is controllable by the agent’s actions, and an exogenous state, which evolves on its own and does not affect the agent’s reward.
-The precise way in which this division exists is unknown to the agent.
-While this may seem similar to our formulation of states and external events, in our case, the external events do influence the dynamics of the state and thereby, the rewards obtained by the agent.
-A similar decomposition of the state space into endogenous and exogenous sub-spaces was studied by Dietterich et al.
-(2018), who assumed that the exogenous states evolve in a Markov fashion and the reward function can be additively decomposed into an exogenous reward and an endogenous reward.
-While their work studies how to decompose the given state into endogenous and exogenous components, in contrast, our work deals with essentially the opposite task, which is to study the effect of incorporating exogenous event information into reinforcement learning algorithms.
-3. Preliminaries
-A Markov Decision Process (MDP) is defined by a tuple $(S,A,\left\{A(s):s\in S\right\},Q,c,\gamma)$, where $S$ is the state space, the set of all possible states of the environment, $A$ is the action space, the set of all possible actions that can be taken by the agent, $A(s)$ is the set of allowed actions at state $s$, $Q$ is the transition probability kernel, $c$ is the cost function, and $\gamma$ is the discount factor.
-At each time step $t$, the agent is in state $s\in S$ and has to take action $a\in A(s)\subseteq A$.
-Depending on the action taken, a cost $c(s,a)$ is incurred by the agent, and the environment transitions to state $s^{\prime}\in S$ sampled from $Q\left(.|(s,a)\right)$, which is then communicated to the agent.
-This process repeats indefinitely.
-In general, $S$ and $A$ are Borel spaces, $A(s)$ is a non-empty measurable subset of $A$ for each $s\in S$, $c$ is a measurable function, and $\gamma\in(0,1)$.
-$Q$ is a stochastic kernel on $S$ given $\mathbb{K}$, where $\mathbb{K}=\left\{(s,a):s\in S,a\in A(s)\right\}$ is the set of all feasible state-action pairs.
-That is, $Q(.|y)$ is a probability measure on $S$ for each $y\in\mathbb{K}$, and $Q(B|.)$ is a measurable function on $\mathbb{K}$ for each $B\in\mathcal{B}(S)$, the Borel $\sigma$-algebra on $S$.
-Any rule $\pi=\{\pi_{t}\}$ by which the agent chooses the action, as a function of state and possibly time, is called a policy.
-A policy can be either deterministic with $\pi_{t}(s)=a_{t}\in A(s)$, or stochastic with $\pi_{t}(.|s\>)$ being a stochastic kernel on $A(s)$ given $s\in S$.
-A policy that is independent of time is called a stationary policy.
-The goal of the agent is to determine a policy $\pi$ among all possible policies $\Pi$ such that the following performance criterion is minimized:
-$V^{\pi}(s)=\mathsf{E}_{s_{t},a_{t}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],\qquad\pi\in\Pi,x\in S,$
-the sum of discounted costs incurred by the agent when it starts from a state and follows policy $\pi$ at each time step, resulting in a trajectory of states $s_{0}=s,s_{1},s_{2},\dots,$ by taking actions $a_{0},a_{1},a_{2},\dots$.
-The function $V^{\pi}$ is called the value function of the policy $\pi$.
-By the Monotone Convergence Theorem, the above can be written as the limit
-$V^{\pi}(s)=\lim_{n\rightarrow\infty}V_{n}^{\pi}(s),$
-where $V_{n}$ is the $n$-stage cost, the sum of discounted costs obtained from the first $n$ time steps.
-A policy $\pi^{*}$ that achieves the minimum value is called an optimal policy and the corresponding value function $V^{\pi^{*}}$ is denoted as $V^{*}$.
-3.1. Bellman optimality equation
-A measurable function $v:S\rightarrow\mathbb{R}$ is said to be a solution to the Bellman optimality equation if it satisfies:
-$v(s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
-The right-hand side of this equation can be considered as an operator $\mathcal{T}$ acting on the function $v$, leading to the optimal Bellman equation
-$v=\mathcal{T}v,$
-where the optimal Bellman operator $\mathcal{T}$ is defined as
-$[\mathcal{T}v](s)=\min_{A(s)}\left[c(s,a)+\alpha\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
-A function that satisfies the optimal Bellman equation is a fixed point of the operator $\mathcal{T}$. Under suitable conditions, the optimal value function satisfies this equation.
-Assumption 1.
-  1. (1)
-    The one-stage cost $c$ is lower semi-continuous, non-negative and inf-compact on $\mathbb{K}$.
-  2. (2)
-    Q is strongly continuous.
-Assumption 2.
-There exists a policy $\pi$ such that $V^{\pi}(s)<\infty$ for all $s\in S$.
-Under the above assumptions, we can make some statements about the existence and uniqueness of the optimal policy and value function, given by the following result.
-Theorem 3.1 (Theorem 4.2.3 of Hernández-Lerma and Lasserre (2012)).
-Suppose the above two assumptions hold, then:
-  1. (1)
-    The optimal value function $V^{*}$ is the point-wise minimal solution to the above Bellman equation. If $u(.)$ is any other solution to the above equation, then $u(.)\geq V^{*}(.)$.
-  2. (2)
-    There exists a function $\pi^{*}:S\rightarrow A$ such that $\pi^{*}(s)\in A(s)$ that attains the minimum in the Bellman equation, i.e.,
-    $V^{*}(s)=c(x,\pi^{*}(s))+\gamma\int V^{*}(s^{\prime})Q(ds^{\prime}|s,\pi^{*}(s))\quad\text{for all }s\in S,$
-    and the deterministic stationary policy $\pi^{*}$ is optimal. Conversely, any deterministic stationary policy that is optimal satisfies the above equation.
-  3. (3)
-    If an optimal policy exists, then there exists one that is deterministic and stationary.
+Organization
+In Section 2, we provide mathematical details of MDPs in a continuous setting to provide a background on conditions for the existence and uniqueness of optimal policy and value function.
+A reader who is familiar with these concepts can skip this Section.
+In Section 3, we set the stage to solve the main problem, and we state the assumptions and regularity conditions.
+We derive guarantees for the existence of almost optimal solutions in Section 4, and a policy iteration algorithm in Section 5.
+We provide sample complexity analysis in Section 6 and experiments in Section 7.
+We provide a detailed literature survey and the position of our work in the existing literature in Section 8.
+2. Preliminaries: Markov Decision Processes
+Consider a Markov Decision Process (MDP) defined by a tuple $\mathcal{M}=(S,A,Q,r,\gamma)$, where $S$ is the state space, the set of all possible states of the environment, $A$ is the action space, the set of all possible actions that can be taken by the agent, $Q$ is the transition probability kernel, $r$ is the reward function, and $\gamma$ is the discount factor.
+At each time step $t$, the agent is in state $s\in S$ and has to take action $a\in A$ obtaining reward $r(s,a)$, and the environment transitions to state $s^{\prime}\sim Q\left(\;.\;|s,a\right)$.
+In general, $S$ and $A$ are Borel spaces [^1] A complete separable metric space equipped with a Borel $\sigma$-algebra, $r$ is a measurable function, and $\gamma\in(0,1)$.
+$Q$ is a stochastic kernel on $S$ given $S\times A$.
+That is, $Q(.|y)$ is a probability measure on $S$ for each $y\in S\times A$, and $Q(B|.)$ is a measurable function on $S\times A$ for each $B\in\mathcal{B}(S)$, the Borel $\sigma$-algebra on $S$.
+A policy is a rule used to choose an action based on the state.
+The goal of the agent is to determine a policy $\pi$ among all possible policies $\Pi$ that achieves the maximum possible value function $V^{\pi}$, which is defined as the expected discounted reward obtained by the agent following policy $\pi$:
+$V^{\pi}(s)=\mathsf{E}_{s_{t},a_{t}}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right],\qquad\pi\in\Pi,s\in S,$
+where the expectation is over the states and actions in the agent’s trajectory due to the stochasticity of the environment and possibly the policy itself.
+A policy $\pi^{*}$ that achieves the optimal value is called an optimal policy, and its corresponding value function $V^{\pi^{*}}$ is denoted as $V^{*}$.
+A measurable function $v:S\rightarrow\mathbb{R}$ is said to be a solution to the Bellman optimality equation if it satisfies $v=\mathcal{T}v$, where $\mathcal{T}$ is the optimal Bellman operator defined as
+$[\mathcal{T}v](s)=\max_{a\in A}\left[r(s,a)+\gamma\int_{S}v(s^{\prime})Q(ds^{\prime}|s,a)\right]\qquad\text{for all }s\in S.$
+A function that satisfies the optimal Bellman equation is a fixed point of the operator $\mathcal{T}$.
+Under suitable regularity conditions on the rewards and transition kernel, the optimal value function $V^{*}$ satisfies this equation.
+Suppose the reward function $r$ is upper semi-continuous, bounded, and sup-compact, and the transition kernel $Q$ is strongly continuous, i.e., the function $v(s,a)=\mathsf{E}_{s^{\prime}\sim Q(.|s,a)}[v(s^{\prime})]$ is continuous and bounded for every measurable bounded function $v$ on $S$.
+Then, there exists a unique optimal policy $\pi^{*}$ whose value function $V^{*}$ is the solution of the Bellman optimality equation (Theorem 4.2.3 of Hernández-Lerma and Lasserre (1996)).
 3. MDP under an External Temporal Process
-Consider an MDP $\mathcal{M}=(S,A,Q,r,\gamma)$ where
-    $S$: The state space, a closed subset of $\mathbb{R}^{m}$
-    $A$: The action space, a closed subset of $\mathbb{R}^{n}$
-    $\mathbb{K}=S\times A$: The set of all feasible state-action pairs
-    $r:S\times A\times S\rightarrow\mathbb{R}$: Reward function as a function of the state-action pair and the resultant next state
-    $\gamma$: Discount factor
-    $Q$: A stochastic kernel on $S$ given $\mathbb{K}$
-We assume that all actions in $A$ are allowed at every state in $S$, so there is no need to specify the allowed action set separately.
-Also, instead of a cost function $c$, we consider a reward function $r$, where the problem objective is maximize the discounted sum of rewards instead.
-Consider an external discrete-time temporal process that is independent of the above MDP but affects the transition probabilities of the MDP.
-The probability of an event occurring in the external process at time $t$ is a function of all the previous (external) events till that time.
+Consider an continuous MDP $\mathcal{M}=(S,A,Q,r,\gamma)$ and an external discrete-time temporal process that is independent of the above MDP but affects the transition probabilities of the MDP.
+The probability of an event occurring in the external process at time $t$ is a function of all the past (external) events till that time.
 Associated with each event is a probabilistic mark $x\in\mathcal{X}$, where $\mathcal{X}$ is a closed subset of $\mathbb{R}^{d}$.
 Suppose at time $t$ that the state of the MDP is $s\in S$.
-Suppose the history of external events till that time is $H_{t}=\{(t_{1},x_{1}),(t_{2},x_{2}),\dots,(t_{j},x_{j})\}$, where the ordered pair $(t_{i},x_{i})$ denotes an event with mark $x_{i}$ that occurred at time $t_{i}$.
-Then, this history perturbs the stochastic kernel $Q$ of the MDP to give rise to a new transition kernel $Q_{H_{t}}$ on $S$ given $\mathbb{K}$.
-Similarly, this history also determines the next event that occurs in the temporal process.
+Denote the history of external events till time $t$ by $H_{t}=\{(t_{1},x_{1}),(t_{2},x_{2}),\dots,(t_{j},x_{j})\}$, where the ordered pair $(t_{i},x_{i})$ denotes an event with mark $x_{i}$ that occurred at time $t_{i}$.
+History $H_{t}$ perturbs the stochastic kernel $Q$ of the MDP to give rise to a new transition kernel $Q_{H_{t}}$ on $S$ given $S\times A$.
+At the same time, this history also determines the next event that occurs in the temporal process.
 We denote the distribution of the event mark at time $t$ by $Q^{X}_{H_{t}}$, which is a probability distribution on $\mathcal{X}$.
 Since this describes the event distribution that is external to the MDP, it does not depend on the state of the MDP or the action taken by the agent.
-4.1. Assumptions
-4.1.1. On the external process
-  1. (i)
-    There exists an event with mark, say $X=0$ (zero), that is equivalent to a non-event.
-  2. (ii)
-    Old events have a reduced influence on the current probability transition kernel of the MDP.
+We assume the following.
+  1. (A1)
+    There exists an event with a mark, say $X=0$ (zero), that is equivalent to a non-event.
+  2. (A2)
+    Events that occurred in the distant past have a reduced influence on the current probability transition kernel of the MDP.
 More specifically, there exists a convergent series $\sum_{T}M_{T}$ of real numbers, such that at any time $t$, for any event histories $H_{t}$ and $H^{\prime}_{t}$, if the MDP is in state $s$ and action $a$ is taken, and if $H_{t}$ and $H^{\prime}_{t}$ differ by only one event at time $t^{\prime}$, then
-    $\mathsf{TV}\left(Q_{H_{t}}(.|s,a),Q_{H^{\prime}_{t}}(.|s,a)\right)<M_{T},\quad\text{for all }(s,a)\in\mathbb{K},$
+    $\mathsf{TV}\left(Q_{H_{t}}(.|s,a),Q_{H^{\prime}_{t}}(.|s,a)\right)<M_{T},\quad\text{for all }s\in S,a\in A,$
     provided $t-t^{\prime}\geq T$. That is, if an event is older than $T$, its influence on the current probability distribution has a Total Variation distance upper bounded by $M_{T}$.
-  3. (iii)
-    Similarly, old events also have a reduced influence on the probability distribution of the current event mark.
+  3. (A3)
+    Similarly, events that occurred in the distant past also have a reduced influence on the probability distribution of the current event mark.
 Specifically, there exists a convergent series $\sum_{T}N_{T}$ of real numbers such that for event histories $H_{t}$ and $H^{\prime}_{t}$ at time $t$ that differ at only one event at time $t^{\prime}$, if $t-t^{\prime}\geq T$, then
     $\mathsf{TV}\left(Q^{X}_{H_{t}},Q^{X}_{H^{\prime}_{t}}\right)<N_{T}.$
-4.1.2. Regularity
-Define a new function $r_{P}:\mathbb{K}\rightarrow\mathbb{R}$ induced from $r$ by $P$ as
+Along with the above assumptions (A1-A3), the MDP $\mathcal{M}$ under the influence of the exogenous event process needs to satisfy some regularity conditions for guaranteeing the existence and uniqueness of optimal solutions.
+Define a new function $r_{P}:S\times A\rightarrow\mathbb{R}$ induced from $r$ by $P$ as
 $r_{P}(s,a)=\mathsf{E}_{s^{\prime}\sim P(.|s,a)}\left[r(s,a,s^{\prime})\right],$
-where $P$ is any stochastic kernel on $S$ given $\mathbb{K}$. We assume the following regularity conditions on $r$ and $Q$:
-  1. (i)
+where $P$ is any stochastic kernel on $S$ given $S\times A$. We assume the following regularity conditions on $r$ and $Q$:
+  1. (B1)
     $r$ is a bounded measurable function. Without any loss in generality, assume its co-domain is $[0,1]\subset\mathbb{R}$.
-  2. (ii)
+  2. (B2)
     The function $r_{P}(s,.)$ is upper semi-continuous for each state $s\in S$ and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process.
-  3. (iii)
-    $r_{P}$ is sup-compact on $\mathbb{K}$, i.e., for every $s\in S,r\in\mathbb{R}$, and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process, the set $\{a\in A:r_{P}(s,a)\geq r\}(\subseteq A)$ is compact.
-  4. (iv)
+  3. (B3)
+    $r_{P}$ is sup-compact on $S\times A$, i.e., for every $s\in S,r\in\mathbb{R}$, and any probability distribution $P=Q_{H}$ induced as a transition kernel by the MDP and any possible external event history $H$ of the temporal process, the set $\{a\in A:r_{P}(s,a)\geq r\}(\subseteq A)$ is compact.
+  4. (B4)
     $Q$ is strongly continuous.
-  5. (v)
+  5. (B5)
     $Q_{H}$ and $Q^{X}_{H}$ are strongly continuous for any feasible event history $H$.
-Note: A transition kernel $P$ is said to be strongly continuous if
+Note that a transition kernel $P$ is said to be strongly continuous if
 $v(s,a)=\mathsf{E}_{s^{\prime}\sim P(.|s,a)}[v(s^{\prime})]$
-is continuous and bounded on $\mathbb{K}$ for every measurable bounded function $v$ on $S$.
+is continuous and bounded on $S\times A$ for every measurable bounded function $v$ on $S$.
 3.1. Examples
 Non-Markov self-exciting event processes that can be described as above are studied extensively in finance, economics, epidemiology, etc., in the form of variants of Hawkes and jump processes.
-While they are studied predominantly in continuous time and sometimes in discrete time, we consider discrete time versions to fit into the traditional discrete-time reinforcement learning framework.
-Consider a discrete-time marked Hawkes process $\left(E_{t},X_{t}\right)_{n\in\mathbb{N}}$ defined as follows:
-$\left(E_{t}\right)_{t\in\mathbb{N}}$ is a sequence of random variables taking values in $\left\{0,1\right\}$, with $E_{t}$ sampled from $\textsf{Bernoulli}\left(p_{t}\right)$, where the intensity $p_{t}$ at time $t$ depends on the realizations of $E_{t}$ at the previous time steps as:
+While they are studied predominantly in continuous time and sometimes in discrete time, we consider discrete-time versions to fit into the traditional discrete-time reinforcement learning framework.
+Consider a discrete-time marked Hawkes process $\left(E_{t},X_{t}\right)_{n\in\mathbb{N}}$ defined as follows: $\left(E_{t}\right)_{t\in\mathbb{N}}$ is a sequence of random variables taking values in $\left\{0,1\right\}$, with $E_{t}$ sampled from $\textsf{Bernoulli}\left(p_{t}\right)$, where the intensity $p_{t}$ at time $t$ depends on the realizations of $E_{t}$ at the previous time steps as:
 $p_{1}$ $=\alpha_{0},$
@@
 The sum $S_{t}=\sum_{t^{\prime}=1}^{t}E_{t}$ is a self-exciting counting process that is the discrete-time counterpart of the continuous-time Hawkes process.
-The above defined sequence of random variables satisfies Assumptions (i) and (ii) required for the external process in Section 4.1.1.
+The above-defined sequence of random variables satisfies Assumptions (A1) and (A2) required for the external process.
 $X_{t}$ is zero when no event occurs, and old events have a reduced influence on current events.
-More precisely, given two histories $(X_{t^{\prime}})_{t^{\prime}<t}$ and $(X^{\prime}_{t^{\prime}})_{t^{\prime}<t}$ that only differ by one event at some time $t^{\prime}\leq t-T$ with $E^{\prime}_{t^{\prime}}=X^{\prime}_{t^{\prime}}=0$ and $E_{t^{\prime}}=1$, $X_{t^{\prime}}=x$, the difference in intensities is $p_{t}-p^{\prime}_{t}=\alpha_{t-t^{\prime}}$, and so,
+More precisely, given two histories $(X_{t^{\prime}})_{t^{\prime}<t}$ and $(X^{\prime}_{t^{\prime}})_{t^{\prime}<t}$ that only differ by one event at some time $t^{\prime}\leq t-T$ with $E^{\prime}_{t^{\prime}}=X^{\prime}_{t^{\prime}}=0$ and $E_{t^{\prime}}=1$, $X_{t^{\prime}}=x$, the difference in intensities is $p_{t}-p^{\prime}_{t}=\alpha_{t-t^{\prime}}$, and hence,
 […]
@@
 $<\alpha_{T}+\textsf{erf}\left(\frac{\beta_{T}}{2\sqrt{2}}\right)=N_{T}.$
-For sufficiently quickly converging $\alpha_{T}$ and $\beta_{T}$, the total variation perturbation $N_{T}$ due to events older than $T$ time steps goes to zero, satisfying Assumption (ii).
+For sufficiently quickly converging $\alpha_{T}$ and $\beta_{T}$, the total variation perturbation $N_{T}$ due to events older than $T$ time steps goes to zero, satisfying Assumption (A2).
 While $(\alpha_{n})$ and $(\beta_{n})$ are any general sequences that need to satisfy some regularity conditions (Seol, 2015) and the above bound, they could also decay in a more structured manner.
-For instance, for some $c>0.5,\lambda>0$ letting
+For instance, for some $c,\lambda>0$ letting
 $\alpha_{n}=ce^{-\lambda n},$
@@
 Similarly, $\beta_{n}$, which determines the distribution of the event, could also be a parametric sequence that decays in an exponential or polynomial fashion.
-5. Characterization of the problem setting
-Given the MDP $(S,A,r,\gamma,Q)$ and an external discrete-time temporal process $\{(t_{i},X_{i})\}_{i}$ that acts on this MDP, we construct a new augmented MDP where the marks of all the events that occurred till time $t$ are appended to the state $s\in S$ to form a new augmented state $\bar{s}\in\bar{S}$, where
-$\bar{s}=(s,x_{t},x_{t-1},x_{t-2},x_{t-3},\dots),\qquad\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}\left(\subset\mathbb{R}^{m}\times\prod_{t=0}^{\infty}\mathbb{R}^{d}\right).$
+The details of notation that is used in the rest of the paper is provided in Table 1.
+4. Guarantees for the existence of almost optimal solutions
+Consider the MDP $\mathcal{M}=(S,A,r,\gamma,Q)$, where $S$ and $A$ are closed subsets of $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, and $r:S\times A\times S\rightarrow\mathbb{R}$ is the reward as a function of the state, action, and resultant next state.
+Given an external discrete-time temporal process $\{(t_{i},X_{i})\}_{i}$ that acts on $\mathcal{M}$, we construct a new augmented MDP where the marks of all the events that occurred till time $t$ are appended to the state $s\in S$ to form a new augmented state $\bar{s}\in\bar{S}$, where
+$\bar{s}=(s,x_{t},x_{t-1},x_{t-2},x_{t-3},\dots),\qquad\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}\left(\subseteq\mathbb{R}^{m}\times\prod_{t=0}^{\infty}\mathbb{R}^{d}\right).$
 Here, $x_{t}$ is the mark of the external event that occurred at time $t$. If no event occurred at time $t$, then $x_{t}=0$.
@@
 Now, the state space of this MDP is infinite-dimensional and poses several challenges.
-However, since we assume that very old events do not affect the transition probabilities during the current time step, we should be able to find a suitable policy that only depends on a finite horizon of past events.
-This is formalized by the following result:
+However, per Assumptions A2 and A3 in Section 3, distant past events do not affect the transition probabilities during the current time step, with the total variation of the effect of events older than $t$ time steps on the transition dynamics and event distribution being bounded by $M_{t}$ and $N_{t}$ respectively.
+This allows one to study the possibility of finding a suitable policy that only depends on a finite horizon of past events.
+The following result formalizes this.
 Theorem 4.1.
-Suppose the MDP and the external process satisfy the assumptions in 4.1.1 and 4.1.2.
+Suppose the MDP and the external process satisfy the assumptions (A1-A3) and the regularity conditions (B1-B5). Then, we have the following.
   1. (1)
@@
     The magnitude of $T$ depends on the rate of convergence of the series $\sum_{t}M_{t}$ and $\sum_{t}N_{t}$.
+This result confirms that the formulated problem is well-posed, with an optimal solution that can be approximated using a tractable policy that is a function of the current state and only a finite history of events.
+This approximation can be as accurate as necessary based on the properties of the exogenous event process and the size of the event history considered.
+The past event horizon $T$ above required for $\epsilon$-suboptimality is some $T$ that satisfies
+$\sum_{t=T+1}^{\infty}M_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}\text{ and}\sum_{t=T+1}^{\infty}N_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}.$
 To prove this, we use a Lemma about the properties of a class of auxiliary MDPs $\mathcal{M}^{(T)}$, whose proof is provided in the Appendix.
-For each time $T>0$, we construct MDP $\mathcal{M}^{(T)}$ as having the same underlying stochastic process as $\mathcal{M}_{X}$, with only one difference: only events in the past $T$ time steps affect the current state transition, and all the events occurring before that have no effect and can effectively be replaced by zeroes.
+For each time $T>0$, we construct MDP $\mathcal{M}^{(T)}_{X}$ as having the same underlying stochastic process as $\mathcal{M}_{X}$, with only one difference: only events in the past $T$ time steps affect the current state transition, and all the events occurring before that have no effect and can effectively be replaced by zeroes.
 Depending on the context, the state space of this MDP can either be taken as $S\times\prod_{t=0}^{\infty}\mathcal{X}$ or $S\times\prod_{t=0}^{T}\mathcal{X}$, since all the older events are zeroed out.
-Any policy $\pi$ for this MDP has an equivalent policy in $\mathcal{M}_{X}$ and vice-versa, where any extra elements in the state are ignored or considered zero.
+Similarly, any policy $\pi$ on the extended state space $\bar{S}$ can operate in both $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$ and thus induces different value functions $V^{\pi}_{\mathcal{M}_{X}}$ and $V^{\pi}_{\mathcal{M}^{(T)}_{X}}$ in these two MDPs respectively.
+Since the MDP $\mathcal{M}_{X}$ is the actual MDP under consideration and $\mathcal{M}^{(T)}_{X}$ is just an approximation, when there is no ambiguity, we use just $V^{\pi}$ without any subscript to denote the value function of policy $\pi$ in the augmented non-stationary MDP $\mathcal{M}_{X}$ and use subscripts when required to distinguish between multiple MDPs.
 Lemma 4.2.
 Let $\pi$ be an arbitrary policy as a function of the augmented state $\bar{s}\in\bar{S}$.
-That is, $\pi$ can be either stochastic or deterministic, and can depend on the current state $s\in S$ and any number of past events.
+That is, $\pi$ can be either stochastic or deterministic and can depend on the current state $s\in S$ and any number of past events in $\mathcal{X}$.
 Then,
-$\left|\left|V^{\pi}_{\mathcal{M}^{(T)}}-V^{\pi}_{\mathcal{M}_{X}}\right|\right|_{\infty}\leq\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{\pi}_{\mathcal{M}^{(T)}}||_{\infty}\right)\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
-where $V^{\pi}_{\mathcal{M}}$ denotes the value function of policy $\pi$ in MDP $\mathcal{M}$.
+$\left|\left|V^{\pi}_{\mathcal{M}^{(T)}_{X}}-V^{\pi}_{\mathcal{M}_{X}}\right|\right|_{\infty}\leq\frac{1}{1-\gamma}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{\pi}_{\mathcal{M}^{(T)}_{X}}\right\|_{\infty}\right)\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
+where $r$ is the reward function and $V^{\pi}_{\mathcal{M}}$ denotes the value function of policy $\pi$ in MDP $\mathcal{M}$.
+Proof.
+See Appendix B.1. ∎
 4.1. Proof of Theorem 4.1
-Part (1) follows from the assumptions and Theorem 3.1.
+Part (1) follows from the assumptions and Theorem 4.2.3 in Hernández-Lerma and Lasserre (1996).
 For part (2), let $\pi^{*}$ be the deterministic optimal stationary policy for $\mathcal{M}_{X}$ and let $\pi^{*(T)}$ be the deterministic optimal stationary policy for $\mathcal{M}^{(T)}$.
-From Lemma 5.2, since $||r||_{\infty}\leq 1$ and $||V||_{\infty}\leq\frac{1}{1-\gamma}$,
-$V^{\pi^{*(T)}}$ $\geq V^{\pi^{*(T)}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
-$\geq V^{\pi^{*}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
+From Lemma 4.2, since $\|r\|_{\infty}\leq 1$ and $\|V\|_{\infty}\leq\frac{1}{1-\gamma}$,
+$V^{\pi^{*(T)}}_{\mathcal{M}_{X}}$ $\geq V^{\pi^{*(T)}}_{\mathcal{M}^{(T)}_{X}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
+$\geq V^{\pi^{*}}_{\mathcal{M}^{(T)}_{X}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
 $\geq V^{\pi^{*}}_{\mathcal{M}_{X}}-\frac{2}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
@@
 $\sum_{t=T+1}^{\infty}M_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}\text{ and}\sum_{t=T+1}^{\infty}N_{t}<\frac{\epsilon(1-\gamma)^{2}}{4}.$
-Choosing such a $T$ gives the Theorem.
+Choosing such a $T$ gives the Theorem. ∎
 5. A policy iteration Algorithm
 Now that an $\epsilon$-optimal policy exists that is a function of only a finite event horizon, we wish to find such a policy.
-We propose a policy iteration algorithm, which alternates between approximate policy evaluation and approximate policy improvement.
-The overall algorithm is given in 1.
+We propose a policy iteration algorithm that alternates between approximate policy evaluation and approximate policy improvement.
+This procedure is listed in Algorithm 1.
 Start with arbitrary deterministic policy $\pi_{0}:S\times\mathcal{X}^{T+1}\rightarrow A$;
@@
 For the policy evaluation step, we only wish to consider candidate value functions that are functions of the finite event horizon, which is why it is not an exact step.
-In this approximate policy evaluation step, we define the value at an infinitely augmented state as a function of the state augmented by its finite event horizon, by sampling the previous older events from some distribution $\mu_{1}$.
-This $\mu_{1}$ is theoretically arbitrary and empirically can be based on the actual event process.
+We define the value function in $\mathcal{M}_{X}$ at an infinitely augmented state as a function of the state augmented by just a finite event horizon by sampling the previous older events from some arbitrary distribution $\mu_{1}$.
+In practice, $\mu_{1}$ is just the actual event process.
 The value function can thus be approximated using Monte Carlo methods.
-For the purpose of analysis, we consider the exact evaluation of the approximate function.
+For the purpose of analysis, we consider the exact evaluation of this approximate value function.
 Similarly, in the policy improvement step, the policy is improved based on the reward and value function resulting from one transition, and this transition can be due to past events coming from any arbitrary distribution $\mu_{2}$.
 The deterministic policy at an augmented state depends only on the finite event horizon and is just the action that maximizes the RHS that depends on the approximate value function.
-Solving the optimization problem in the policy improvement step requires knowledge of the models $r(s,a,s^{\prime})$, $Q$ and $Q^{X}$.
-In this work, we ignore any approximation errors arising due to estimating the models and analyze the algorithm assuming the exact step is taken as defined.
+Solving the optimization problem in the policy improvement step requires knowledge of the models $r(s,a,s^{\prime})$, $Q$, and $Q^{X}$.
+In this work, we ignore any approximation errors arising due to estimating the models and analyze the algorithm by assuming the exact step is taken as defined.
 5.1. Analysis
-For the analysis of this algorithm, we need the following Lemma, which bounds the difference in the value of the value function at two different augmented states when they differ in events that are more than $T$ time steps old.
+For the analysis of this algorithm, we need the following Lemma, which bounds the difference in the value function at two different augmented states when they differ in events that are more than $T$ time steps old.
 Lemma 5.1.
@@
 $\sup_{\bar{s}=(s,x_{0:\infty})}\left|V^{\pi}(\bar{s})-V^{\pi}((s,x_{0:T},(0)_{T+1:\infty}))\right|<\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
+Proof.
+See Appendix B.2. ∎
+Here, $(s,x_{0:T},(0)_{T+1:\infty})=(s,x_{0},x_{1},\dots,x_{T},0,0,\dots)\in\bar{S}$ is the extended state that has been essentially “truncated” and made finite by replacing events older than $T$ time steps by zero.
 Corollary 5.2.
 For a given policy $\pi$ that is a function of events only in the past $T$ time steps, and for any two augmented states $\bar{s}_{1}$ and $\bar{s}_{2}$ that differ in any number of events that occurred before the previous $T$ time steps,
-$\left|V^{\pi}(\bar{s}_{1})-V^{\pi}(\bar{s}_{2})\right|\leq\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$ (1)
-Throughout the rest of this section, we let $\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)$, the change in the value function induced by changing an arbitrary number of sufficiently old events.
-Now, we characterize the behavior of the policy iteration procedure described previously.
-In general, exact policy iteration results in a sequence of policies that converge to the optimal policy.
-However, our approximate policy iteration produces policies whose value functions satisfy the following Lemma.
+$\left|V^{\pi}(\bar{s}_{1})-V^{\pi}(\bar{s}_{2})\right|\leq\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
+Essentially, this means that if a policy considers only events in the past $T$ time steps, its value at states differing at older events differ by an amount proportional to the extent of non-stationarity induced by all events older than $T$ steps.
+This helps characterize the behavior of the policy iteration procedure described previously.
+While exact policy iteration results in a sequence of policies that converge to the optimal policy, our approximate policy iteration produces policies whose value functions satisfy the following Lemma, whose proof is given later in this section.
 Lemma 5.3.
 $V^{\pi_{k+1}}\geq V^{\pi_{k}}-\frac{2+5\gamma}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t}).$
-That is, there is no guarantee that the sequence of value functions is non-decreasing.
-The only guarantee is that they do not decrease by more than the specific amount given by the Lemma.
-This is due to the approximation error $\epsilon$ at each time step.
+That is, while there is no guarantee that the sequence of value functions $\left\{V^{\pi_{k}}\right\}$ is non-decreasing, Lemma 5.3 provides a guarantee that they do not decrease by more than a specific amount.
+This is due to the approximation error
+$\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
+at each time step induced by the non-stationarity due to the exogenous events.
 In general, approximate policy iteration algorithms converge under the condition that the approximation errors slowly vanish over the course of the algorithm.
 However, our approximation error of $\epsilon$ remains constant throughout, leading to the above situation.
-However, intuitively it would seem that such a small approximation error $\epsilon$ should only disturb the convergence of the algorithm when we are $\epsilon$-close to the optimal solution.
-During the initial iterations when we are very far from the solution, such small approximation errors should not matter.
-It turns out that this is actually the case in some sense.
-Degradation of the value function occurs only in parts of the state space where the Bellman error is close to zero.
+Intuitively it would seem that such a small approximation error $\epsilon$ should only disturb the convergence of the algorithm when we are $\epsilon$-close to the optimal solution.
+During the initial iterations, when we are very far from the solution, such small approximation errors should not matter.
+We next establish theoretically that the degradation of the value function occurs only in parts of the state space where the Bellman error is close to zero.
 A Bellman error of zero generally, although not always, corresponds to the optimal policy.
@@
 Theorem 5.4.
-At iteration $k$, for any augmented state $\bar{s}\in S\times\prod_{t=0}^{\infty}\mathcal{X}$, at least one of the following holds:
+At iteration $k$ of the policy iteration procedure given in Algorithm 1, for any augmented state $\bar{s}\in\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}$, at least one of the following holds:
   1. (1)
-    $V^{\pi_{k+1}}(\bar{s})\geq V^{\pi_{k}}(\bar{s})$.
+    $V^{\pi_{k+1}}(\bar{s})\geq V^{\pi_{k}}(\bar{s})$,
   2. (2)
-    $|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})|<\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$.
+    $|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})|<\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$,
+where $V^{\pi}$ denotes the value function of policy $\pi$ in the MDP $\mathcal{M}_{X}$.
 That is, the performance of the policy improves everywhere except the region of the state space where the Bellman error is very small.
-The faster the decay of exogenous event influence, the lesser the approximation error and smaller the region of state space where the policy may not improve.
+The faster the decay of exogenous event influence, the lesser the approximation error and the smaller the region of state space where the policy may not improve.
 Proof.
@@
 $\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\\ a=\pi_{k}(\bar{s})\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]$
-$<\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]-\delta$
+$\qquad\qquad\qquad<\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]-\delta$
 for some suitable $\delta>0$ yet to be chosen.
 It is to be noted that whether or not this condition holds, the above inequality or the corresponding equality holds for $\delta=0$ by the definition of $\pi_{k+1}$.
-Suppose the above condition holds. Then,
+Suppose the above condition holds. Then, for $\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$,
 $V^{\pi_{k}}((s,x_{0:\infty}))$ $\leq\mathsf{E}_{x^{\prime}_{T+1:\infty}\sim\mu_{2}}V^{\pi_{k}}((s,x_{0:T},x^{\prime}_{T+1:\infty}))+\epsilon$
@@
 $=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:\infty}}(.)\\ a=\pi_{k+1}(\bar{s})\end{subarray}}r(s,a,s^{\prime})+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:\infty}}(.)\\ a=\pi_{k+1}(\bar{s})\end{subarray}}V^{\pi_{k}}((s^{\prime},x^{\prime},x_{0:\infty}))$
-$\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\gamma\epsilon+\epsilon-\delta$
+$\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\gamma\epsilon+\epsilon-\delta.$
 That is, $V^{\pi_{k}}(\bar{s})$ can be bounded recursively in terms of the reward obtained using $\pi_{k+1}$ and $V^{\pi_{k}}(\bar{s}^{\prime})$, where $\bar{s}^{\prime}$ comes from following policy $\pi_{k+1}$.
@@
 $V^{\pi_{k}}(\bar{s})$ $\leq\mathsf{E}_{\pi_{k+1}}\sum_{t=0}^{\infty}\gamma^{t}r(s,a,s^{\prime})-\delta$
-$\qquad\qquad+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
+$\qquad\qquad\qquad\qquad+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
 $=V^{\pi_{k+1}}(\bar{s}),$
@@
 From this, it can be seen that as the time window $T$ considered increases, the error reduces as $e^{-\left\{\bar{\lambda},\lambda_{\alpha},\lambda_{\beta}\right\}T}$.
-So, depending on the rate of influence decay, increasing the time window beyond a certain point gives diminishing returns.
+Hence, depending on the rate of influence decay, increasing the time window beyond a certain point gives diminishing returns.
 The error also has a proportional dependence on $c$, the base intensity of the Hawkes process and an inverse dependency on $\lambda_{\alpha}$, the decay of the excitation function of the Hawkes process.
 6. Sample Complexity
-The algorithm given in Section 6 defines the value function and its update.
-But in practice, the value function has to be learnt using samples obtained from the environment.
-Further, the actual value function is generally approximated from a chosen set of functions.
-In this section, we analyze the sample complexity of policy evaluation by using samples from a single sample path induced by the policy, and give error bounds on the evaluation error for the case of a linear function space and some regularity assumptions on the MDP.
+While the algorithm given in Section 5 defines the value function and its update, in practice, it has to be learnt using samples obtained from the environment and is generally approximated from a chosen class of functions.
+In this section, we analyze the sample complexity of policy evaluation by pathwise Least-Squares Temporal Difference (LSTD) (Lazaric et al., 2012), which uses samples from a single sample path induced by the policy, and give error bounds on the evaluation error for the case of a linear function space and some regularity assumptions on the MDP.
+In contrast to Lazaric et al.
+(2012), our work considers an MDP with an infinite-dimensional augmented state space, a stochastic reward function and an additional error due to the use of tractable policies and value functions on a finite-dimensional domain.
+This results in additional sample complexity even for evaluating a given policy, depending on the extent to which the exogenous events affect the evolution of the states.
+In this section, we quantify the precise nature of this error.
 Function approximation
@@
 Since the domain of the functions in this class is infinite-dimensional, this can be further broken down into two approximations for dealing with these functions practically: a truncation operation to make the domain finite dimensional, and then another class of functions that are then used to cover possible functions on this domain, i.e,
-=$\mathcal{F}=\left\{f\;o\;f^{T}_{\text{trunc}}:f\in\mathcal{F}^{T}\right\},$
-where $f_{\text{trunc}}:\bar{S}\rightarrow S\times\bigcup_{t=0}^{T}\mathcal{X}$ removes extra events that are older than $T$ time steps while preserving the current state and new events, and $\mathcal{F}^{T}$ is some function class of functions defined on this new truncated domain.
-For a fixed $f^{T}_{\text{trunc}}$, finding the best function in $\mathcal{F}$ is the same as finding the best function in $\mathcal{F}^{T}$.
+=$\mathcal{F}=\left\{f\;o\;f^{(T)}_{\text{trunc}}:f\in\mathcal{F}^{(T)}\right\},$
+where $f^{(T)}_{\text{trunc}}:\bar{S}\rightarrow S\times\bigcup_{t=0}^{T}\mathcal{X}$ removes extra events that are older than $T$ time steps while preserving the current state and new events, and $\mathcal{F}^{(T)}$ is some function class of functions defined on this new truncated domain.
+For a fixed $T$, finding the best function in $\mathcal{F}$ is the same as finding the best function in $\mathcal{F}^{(T)}$.
 Finally, we wish to break down the error between the true value function and the learned truncated value function in terms of the errors introduced due to these two operations as well as the stochasticity of the samples used to learn the approximate function.
 We wish to bound
-$\left|\left|V-\tilde{V}\right|\right|_{\rho},$
-where $V$ is the true value function on $\bar{S}$, $\tilde{V}$ is the learned value function in $\mathcal{F}$, and $\rho$ is some distribution over $\bar{S}$, and the norm is the $l^{2}(\rho)$-norm, which is the expected $l^{2}$ norm of the value of the function w.r.t the measure $\rho$, i.e,
-$\left|\left|f\right|\right|^{2}_{\rho}=\int_{\bar{S}}f(x)^{2}\;\rho(\mathrm{d}x).$
-While the $f^{T}_{trunc}$ operator just ignores old events, we also define a projection operator $\Pi^{trunc}$ onto the function space defined on the truncated domain, as
-$\Pi^{trunc}V=\inf_{f:S\times\bigcup_{t=0}^{T}\mathcal{X}}\left|\left|f-V\right|\right|_{\rho}.$
+$\left\|V-\widetilde{V}\right\|_{\rho},$
+where $V$ is the true value function on $\bar{S}$, $\widetilde{V}$ is the learned value function in $\mathcal{F}$, and $\rho$ is some distribution over $\bar{S}$, and the norm is the $l^{2}(\rho)$-norm, which is the expected $l^{2}$ norm of the value of the function w.r.t the measure $\rho$, i.e,
+$\left\|f\right\|^{2}_{\rho}=\int_{\bar{S}}f(x)^{2}\;\rho(\mathrm{d}x).$
+While the $f^{(T)}_{trunc}$ operator just ignores old events, we also define a projection operator $\Pi^{trunc}$ onto the function space defined on the truncated domain, as
+$\Pi^{trunc}V=\inf_{f:S\times\bigcup_{t=0}^{T}\mathcal{X}\rightarrow\mathbb{R}}\left\|f-V\right\|_{\rho}.$
 which finds the best approximation that does not depend on the old events, where ‘best’ is defined in expectation with respect to $\rho$.
@@
 A standard approximating function space considered is a an arbitrary finite dimensional function space.
-Specifically, assume that there are $d$ basis functions $\varphi_{i}:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}$ that linearly span $\mathcal{F}$ and are bounded by $L$, that is, for each $f\in\mathcal{F}$, there exist $\alpha\in\mathbb{R}^{d}$ such that $f=\sum_{i\in[d]}\alpha_{i}\varphi_{i}$.
+Specifically, assume that there are $d$ basis functions $\varphi_{i}:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}$ that linearly span $\mathcal{F}^{(T)}$ and are bounded by $L$, that is, for each $f\in\mathcal{F}^{(T)}$, there exist $\alpha\in\mathbb{R}^{d}$ such that $f=\sum_{i\in[d]}\alpha_{i}\varphi_{i}$.
 These basis functions together form a feature representation function $\phi:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}^{d}$, defined by $\phi(x)=(\varphi_{1}(x),\dots,\varphi_{d}(x))$.
 6.1. Approximate policy evaluation
-The best approximate value function $\tilde{V}$ is that function in $\mathcal{F}$ (or equivalently $\mathcal{F}^{T}$) that minimizes the above expected norm difference.
+The best approximate value function $\widetilde{V}$ is that function in $\mathcal{F}$ (or equivalently $\mathcal{F}^{(T)}$) that minimizes the above expected norm difference.
 Suppose we are given just $N$ samples $\bar{s}_{1},\dots,\bar{s}_{N}$ from a Markov chain induced by policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$, and corresponding rewards $r_{1},\dots,r_{N}$.
-The next best thing is to minimize the empirical norm difference at these points, defined as
+We minimize the empirical norm difference at these points, defined as
 $\left|\left|f\right|\right|_{N}=\left(\frac{1}{N}\sum_{t=1}^{N}f(x_{t})^{2}\right)^{\frac{1}{2}}.$
@@
 $\mathcal{F}_{N}$ =$=\left\{\left(f(\bar{s}_{1}),\dots,f(\bar{s}_{N})\right):f\in\mathcal{F}\right\}$
-=$=\left\{\left(f(s^{T}_{1}),\dots,f(s^{T}_{N})\right):f\in\mathcal{F}^{T}\right\}$
+=$=\left\{\left(f(s^{T}_{1}),\dots,f(s^{T}_{N})\right):f\in\mathcal{F}^{(T)}\right\}$
 =$=\left\{\Phi\alpha:\alpha\in\mathbb{R}^{d}\right\},$
-where $\Phi=\left(\phi(\bar{s}_{1}^{T}),\dots,\phi(\bar{s}_{N}^{T})\right)_{N\times d}$, and $\bar{s}_{i}^{T}=f^{T}_{trunc}(\bar{s})$.
+where $\Phi=\left(\phi(\bar{s}_{1}^{T}),\dots,\phi(\bar{s}_{N}^{T})\right)_{N\times d}$ and $\bar{s}_{i}^{T}=f^{(T)}_{trunc}(\bar{s})$.
 The true value function $V$ is a fixed point of the Bellman operator $\mathcal{T}$ in $\bar{S}$.
-From the given $N$ points, we can define our approximate value function as the fixed point of the projected pathwise Bellman operator $\hat{\Pi}\hat{\mathcal{T}}$, where $\hat{T}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the pathwise Bellman operator, with $\hat{T}(x)$ being defined for $x\in\mathbb{R}^{N}$ as
-$\left(\hat{T}(x)\right)_{i}=\begin{cases}r_{i}+\gamma x_{i+1}&\text{ for }i<N\\ r_{i}&\text{ for }i=N\end{cases},$
+From the given $N$ points, we can define our approximate value function as the fixed point of the projected pathwise Bellman operator $\hat{\Pi}\hat{\mathcal{T}}$, where $\hat{\mathcal{T}}:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ is the pathwise Bellman operator, with $\hat{\mathcal{T}}(x)$ being defined for $x\in\mathbb{R}^{N}$ as
+$\left(\hat{\mathcal{T}}(x)\right)_{i}=\begin{cases}r_{i}+\gamma x_{i+1}&\text{for }i<N\\ r_{i}&\text{ for }i=N\end{cases}$
 and $\hat{\Pi}:\mathbb{R}^{N}\rightarrow\mathcal{F}_{N}$ is the projection operator onto the valid function approximated values on $x_{1:N}$, i.e,
-$\hat{\Pi}(x)=\mathop{\mathrm{argmin}}\limits_{z\in\mathcal{F}_{N}}\left|\left|x-z\right|\right|_{N}^{2}.$
-This also be written as $\hat{T}=r+\gamma\hat{P}$, where $r=(r_{t})_{t\in[N]}$ and $\hat{P}$ is the path-wise transition operator.
-$\hat{T}$ is a $\gamma$-contraction in ||.$||.||_{N}$ as well as ||.$||.||_{2}$, and the projection operator $\hat{\Pi}$ onto $\mathcal{F}_{N}$ is non-expansive since $\mathcal{F}$ is convex.
-This means that $\hat{\Pi}\hat{T}$ is a $\gamma$-contraction on $\mathbb{R}^{N}$ and has a unique fixed point $\hat{v}\in\mathcal{F}_{N}$, which we define as our approximation of the value function values at these points.
+$\hat{\Pi}(x)=\mathop{\mathrm{argmin}}\limits_{z\in\mathcal{F}_{N}}\left\|x-z\right\|_{N}^{2}.$
+This also be written as $\hat{\mathcal{T}}=r+\gamma\hat{P}$, where $r=(r_{t})_{t\in[N]}$ and $\hat{P}$ is the path-wise transition operator.
+$\hat{\mathcal{T}}$ is a $\gamma$-contraction in .$\|.\|_{N}$ as well as .$\|.\|_{2}$, and the projection operator $\hat{\Pi}$ onto $\mathcal{F}_{N}$ is non-expansive since $\mathcal{F}$ is convex.
+This means that $\hat{\Pi}\hat{\mathcal{T}}$ is a $\gamma$-contraction on $\mathbb{R}^{N}$ and has a unique fixed point $\hat{v}\in\mathcal{F}_{N}$, which we define as our approximation of the value function values at these points.
 Corresponding to this point, there exists function $\hat{V}\in\mathcal{F}$ such that $\hat{v}=\left(\hat{V}(\hat{s}_{t})\right)_{t\in[N]}$.
@@
 This $\hat{V}$ is considered the output of the approximate policy evaluation step of the policy iteration algorithm.
-For determining the overall error of this estimate, the intermediate task is to bound the difference between this and the true solution on these data points, i.e, $||V-\hat{V}||_{N}$, as a function of the number of data points and the complexity of the function class under consideration.
+For determining the overall error of this estimate, the intermediate task is to bound the difference between this and the true solution on these data points, i.e, $\|V-\hat{V}\|_{N}$, as a function of the number of data points and the complexity of the function class under consideration.
 Let $v=\left(V(\bar{s}_{t})\right)_{t\in[N]}$ just as $\hat{v}=\left(\hat{V}(\bar{s}_{t})\right)_{t\in[N]}$. Following Lazaric et al. (2012),
-$||V-\hat{V}||_{N}=||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}.$
+$\left\|V-\hat{V}\right\|_{N}=\left\|v-\hat{v}\right\|_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{1}{1-\gamma}\left\|\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v\right\|_{N}.$
 The first term is the approximation error due to having a restricted function space and is unavoidable.
 The second term also includes the estimation error due to using the pathwise Bellman operator instead of the actual Bellman operator, and is a function of the number of samples used.
-It can be written as $||\hat{\Pi}\xi||_{N}$, where $\xi=\hat{\mathcal{T}}v-v$ is the estimation error.
+It can be written as $\left\|\hat{\Pi}\xi\right\|_{N}$, where $\xi=\hat{\mathcal{T}}v-v$ is the estimation error.
 Now, for $t<N$,
@@
 $\qquad-\gamma\mathsf{E}_{\bar{S}_{t+1}\sim Q,Q^{X}_{x_{:t}}(.|s_{t},\pi(\bar{s}_{t}))}\left[V(\bar{S}_{t+1})\right].$
-$\xi_{t}$ are functions of $\bar{s}_{t}$, that are samples from the Markov chain induced by the policy $\pi$ in the augmented MDP $\bar{M}$.
+$\xi_{t}$ are functions of $\bar{s}_{t}$, that are samples from the Markov chain induced by the policy $\pi$ in the augmented MDP $\mathcal{M}_{X}$.
 Considering them as a random variables, $\left(\xi_{t}\varphi_{t}(\bar{s_{t}})\right)_{1\leq t<N}$ is a Martingale difference sequence w.r.t $\bar{s}_{t}$, with each element being zero mean and bounded by $\frac{L}{1-\gamma}$.
-So, applying Azuma’s inequality to $(\xi_{t}\varphi_{i}(\bar{s})_{t})_{1\leq t<N}$ gives, for each $i\in[d]$,
+So, applying Azuma’s inequality to $(\xi_{t}\varphi_{i}(\bar{s}_{t}))_{1\leq t<N}$ gives, for each $i\in[d]$,
 $\mathsf{P}\left(\left|\sum_{t=1}^{N-1}\xi_{t}\varphi_{i}(s_{t})\right|\geq\epsilon\right)<2\exp\left(-\frac{\epsilon^{2}(1-\gamma)^{2}}{2(N-1)L^{2}}\right).$
@@
 $\left|\sum_{t=1}^{N-1}\xi_{t}\varphi_{i}(s_{t})\right|<\frac{L}{1-\gamma}\left(\sqrt{2(N-1)\log(2d/\delta)}+1\right)\;\forall i\in[d]\text{ w.p. }\geq 1-\delta.$
-Following Lazaric et al. (2012), this can be used to bound $||\hat{\Pi}\xi_{n}||_{N}$ as
-$\left|\left|\hat{\Pi}\hat{\mathcal{T}}v-\hat{\Pi}v\right|\right|_{N}\leq\frac{1}{N}\sqrt{\frac{d}{\nu_{n}}}\max_{i}\left|\sum_{t=1}^{N}\xi_{t}\varphi_{i}(s_{t})\right|,$
+Following Lazaric et al. (2012), this can be used to bound $\|\hat{\Pi}\xi_{n}\|_{N}$ as
+$\left\|\hat{\Pi}\hat{\mathcal{T}}v-\hat{\Pi}v\right\|_{N}\leq\frac{1}{N}\sqrt{\frac{d}{\nu_{n}}}\max_{i}\left|\sum_{t=1}^{N}\xi_{t}\varphi_{i}(s_{t})\right|,$
 where $\nu_{n}$ is the smallest positive eigenvalue of the Gram matrix $\Phi^{T}\Phi$.
 Substituting these terms in the original inequality, we obtain
-$||v-\hat{v}||_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{1}{1-\gamma}||\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v||_{N}$
-$\leq\frac{1}{\sqrt{1-\gamma^{2}}}||v-\hat{\Pi}v||_{N}+\frac{L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu_{N}}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right).$ (2)
+$\left\|v-\hat{v}\right\|_{N}$ $\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{1}{1-\gamma}\left\|\hat{\Pi}v-\hat{\Pi}\hat{\mathcal{T}}v\right\|_{N}$
+$\leq\frac{1}{\sqrt{1-\gamma^{2}}}\left\|v-\hat{\Pi}v\right\|_{N}+\frac{L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu_{N}}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right).$ (1)
 The above identity is almost identical to the one in Lazaric et al.
 (2012), the reason being that we are still operating in the range space of the basis functions restricted to a finite set of states and so the structure of the MDP in our setting does not affect the analysis.
-Since the proof follows a similar path, we have only included those steps where there is a discrepancy, which are mainly due to the noise terms taken in the martingale difference sequence.
+Since the proof follows a similar path, we have only included those steps that differ from the original proof, which are mainly due to the noise terms taken in the martingale difference sequence.
 In our setting, these noise terms also include stochasticity due to the reward in addition to the transition function, leading to a difference in the final expression.
-While the final inequality is quite similar, we shall see that when considering the expected error in the augmented state space, we can bring out the properties of our setting by splitting the expected error in terms of error due to function truncation and the inherent error due to linear function approximation.
+While this final inequality is quite similar, we shall see that when considering the expected error in the augmented state space, we can bring out the properties of our setting by splitting the expected error in terms of error due to function truncation and the inherent error due to linear function approximation.
 6.2. Expected error
@@
 Theorem 6.1.
-Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using policy $\pi$ in the MDP $\mathcal{M}_{X}$.
-Suppose sufficiently large $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$.
-Let $\nu$ be a lower bound on the eigenvalues of the sample-based Gram matrix that hold with probability $41-\delta/$ 4.
+Let $\bar{s}_{1},\dots,\bar{s}_{N+\widetilde{N}}$ be a sample path obtained using policy $\pi$ in the MDP $\mathcal{M}_{X}$.
+Suppose sufficiently large $\widetilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ samples are used to compute the truncated least-squares path-wise estimate $\widetilde{V}$ of the true value function $V$.
+Let $\nu$ be a lower bound on the eigenvalues of the sample-based Gram matrix that holds with probability $41-\delta/$ 4.
 Then, with probability at least $1-\delta$,
-$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
-$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(8d/\delta)}{N}}+\frac{1}{N}\right)+\bar{\epsilon}(N,d,\delta).$
-Proof.
-Following proof of Theorem $5$ of Lazaric et al. (2012),
-$2||\hat{V}-V||_{N}\geq 2||\tilde{V}-V||_{N}\geq||\tilde{V}-V||_{\rho}-\epsilon_{1},$
-where the first inequality is because of truncation, and the second inequality is a generalization Lemma for Markov chains, and so
-$||\tilde{V}-V||_{\rho}$ $\leq 2||\hat{V}-V||_{N}+\epsilon_{1}$
-$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\hat{\Pi}V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
-$\leq\frac{2}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
-$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}||V-\Pi V||_{\rho}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
-$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-\Pi^{trunc}V||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
+$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
+$\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(8d/\delta)}{N}}+\frac{1}{N}\right)+\bar{\epsilon}(N,d,\delta),$
+where $V$ is the true value function, $\widetilde{V}$ is the learned value function clipped at $\frac{1}{1-\gamma}$, $\Pi^{trunc}V$ is the best possible value function on the truncated state space, $\Pi V$ is the best approximating value function in the linear function space considered, $M_{t}$ and $N_{t}$ are the upper bounds on the total variation due to exogenous events older than $t$ time steps induced in the transition dynamics and event mark distribution respectively, and $\bar{\epsilon}$ depends on the mixing rate of the Markov chain, dimensionality $d$ of the features, and the probability of failure $\delta$.
+This result breaks down the overall error into individual components, viz.
+the approximation error due to the non-stationarity, the inherent error due to linear function approximation, the error due to stochasticity of the samples that reduces with the number of samples, and other errors due to mixing of the Markov chain, the dimensionality of the function space, etc.
+The detailed conditions on the MDP under which this result holds, as well as the exact expression for $\bar{\epsilon}$, are given in the Appendix.
+Proof of Theorem 6.1.
+Following proof of Theorem $5$ in Lazaric et al. (2012),
+$2\left\|\hat{V}-V\right\|_{N}\geq 2\left\|\widetilde{V}-V\right\|_{N}\geq\left\|\widetilde{V}-V\right\|_{\rho}-\epsilon_{1},$
+where the first inequality results from truncation, and the second inequality is a generalization Lemma for Markov chains, and so
+$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq 2\left\|\hat{V}-V\right\|_{N}+\epsilon_{1}$
+$\leq\frac{2}{\sqrt{1-\gamma^{2}}}\left\|V-\hat{\Pi}V\right\|_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
+$\leq\frac{2}{\sqrt{1-\gamma^{2}}}\left\|V-\Pi V\right\|_{N}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}$
+$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left\|V-\Pi V\right\|_{\rho}+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
+$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-\Pi^{trunc}V\right\|_{\rho}+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
 $\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
-$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
+$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-V^{\epsilon}(s,x_{0:T},\bar{0})\right\|_{\rho}+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
 $\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
-$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(||V-V^{\epsilon}||_{\rho}+||V^{\epsilon}-V^{\epsilon}(s,x_{0:T},\bar{0})||_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
+$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\left\|V-V^{\epsilon}\right\|_{\rho}+\left\|V^{\epsilon}-V^{\epsilon}(s,x_{0:T},\bar{0})\right\|_{\rho}+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
 $\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}$
-$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
+$\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
 $\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta^{\prime})}{N}}+\frac{1}{N}\right)+\epsilon_{1}+2\sqrt{2}\epsilon_{2}.$
@@
 (2002), require the domain of the basis functions to be finite dimensional, the analysis takes place in the co-domain and so can be directly applied to our setting.
-We state the complete theorem along with the Lemmas in Appendix B.
+We state the complete theorem along with the Lemmas in Appendix C.
 7. Experiments
-We conducted experiments for solving a non-stationary variant of classic Pendulum-v1 environment in the control suite of OpenAI Gym.
+We conducted experiments for analyzing policy evaluation in a non-stationary variant of the classic Pendulum-v1 environment in the control suite of OpenAI Gym.
 The task consists of applying torque to a pendulum to swing it and keep it upright.
-We considered a fixed neural-network policy that has been pre-trained partially using DDPG.
+We considered a fixed neural-network policy that has been pre-trained partially using DDPG (Lillicrap et al., 2016).
 We then used pathwise LSTD with linear function approximation for evaluating this policy as described in Section 6.
@@
 This is because the events older than the horizon now contribute a negligible influence on the current dynamics, while the dimensionality of the features is increasing due to the inclusion of more past events in the features, resulting in a slightly greater expected error.
-The third plot shows the dependence on $\lambda_{\alpha}$, the rate of decay of $\alpha_{t}$.
+The first plot shows the dependence on $\lambda_{\alpha}$, the rate of decay of $\alpha_{t}$.
 Faster decay results in lesser non-stationarity and lower approximation error due to the first term in the sample complexity upper bound.
@@
 The error bars indicate the range of $40$-$60$ percentile versus the decay rate and $20$-$80$ percentile for the other two.
+These results empirically demonstrate the sample complexity bound for policy evaluation with linear function approximation in Theorem 6.1.
+The average error decreases with increasing $N$ and decreasing $N_{t}$.
+Increasing the event horizon $T$ increases $d$ while reducing $\sum_{t=T+1}^{\infty}\{M_{t},N_{t}\}$, resulting in a trade-off between the corresponding two terms in the bound, with the minimal error being achieved for $T=2$.
+8. Related Work
+Many previous studies in the literature that address non-stationarity primarily provide practical solutions that lack theoretical backing.
+While some research does offer theoretical insights in specific special cases, our work tackles this problem from a comprehensive perspective.
+The nature of the results reported in this paper is significantly different from those available in the existing literature.
+8.1. Practical approaches in the literature
+Numerous previous studies examine environments characterized by piece-wise stationarity, where arbitrary changes transpire at designated change points.
+Different studies adopt differing assumptions regarding the prior knowledge accessible to the agent concerning the nature of these change points.
+For instance, Padakandla et al.
+(2020) assume that the transition and reward functions belong to a fixed set of ‘contexts’ between change points.
+This work introduces a model-free algorithm that employs a change point detection mechanism to ascertain the timing changes by analyzing the state-reward-next state tuples obtained from the environment.
+Although this can be regarded as a special case within our framework, it is important to note that the context remains unobserved in this scenario, while the pattern of change in the context is presumed to be known a priori.
+Further, they consider only a fixed set of potential contexts.
+Due to these issues, the proposed algorithm in Padakandla et al.
+(2020) cannot be applied in our case since we consider a continuous space of potential contexts and no knowledge of the order in which the contexts will occur.
+Alegre et al.
+(2021) proposed a model-based algorithm that can deal with arbitrarily different types of contexts and continuous state and action spaces.
+Although this approach considers a relatively more general setting, the focus of this work is to detect the change points efficiently and to learn separate models for each context.
+This is infeasible for our setting where the context changes possibly every time step.
+Hadoux et al.
+(2014) formalized a class of piece-wise stationary MDPs called Hidden-Semi-Markov-Model MDPs where the transition from one hidden stationary MDP mode to another is semi-Markov in nature.
+To solve this problem, this work adapts the Partially Observable Monte Carlo Planning (POMCP) algorithm used for solving Partially Observable Markov Decision Processes (POMDPs).
+This work, too, does not provide any theoretical results.
+8.2. Theoretical results available in the literature
+Lecarpentier and Rachelson (2019) considers a finite horizon MDP in which the transition dynamics and reward function evolve such that they are Lipschitz continuous with respect to time.
+Cheung et al.
+(2020) considers such a setting in a finite MDP where the amount of total change in the environment, referred to as the variation budget, is known to the agent.
+This work proposes an optimistic Value Iteration based algorithm and provides guarantees on the upper bound of its dynamic regret.
+Wei and Luo (2021) proposes a black-box reduction that turns any suitable algorithm with optimal performance in a near-stationary environment into an algorithm that achieves optimal dynamic regret in a non-stationary environment.
+Some algorithms assume additional structure on the nature of non-stationarity.
+Feng et al.
+(2022) propose Factored Non-stationary MDPs (FN-MDPs), where the transition dynamics are defined in terms of a Dynamic Bayesian Network over the states, actions, rewards, and latent change factors that cause non-stationarity.
+Here, both the states and the latent change factors have transition dynamics obeying this causal graph, and the latent factors evolve in a Markov fashion.
+They propose using a Variational Auto-Encoder (VAE) framework to learn transition dynamics and present results in MuJoCo and robotic manipulation tasks.
+A factorization setting with a resemblance to ours is the Exogenous MDP (Efroni et al., 2022), wherein the environment state is divided into two parts, an endogenous state which is controllable by the agent’s actions, and an exogenous state, which evolves on its own and does not affect the agent’s reward.
+The precise way in which this division exists is unknown to the agent.
+While this may seem similar to our formulation of states and external events, in our case, the external events do influence the dynamics of the state and, thereby, the rewards obtained by the agent.
+Further, Efroni et al.
+(2022) only consider the tabular setting and propose algorithms that can only deal with finite horizon episodes.
+A similar decomposition of the state space into endogenous and exogenous sub-spaces was studied in Dietterich et al.
+(2018), where it is assumed that the exogenous states evolve in a Markov fashion and the reward function can be additively decomposed into an exogenous reward and an endogenous reward.
+This work proposed algorithms to determine the exogenous component of the state as projections that maximize the partial correlation coefficient as a proxy for the mutual information between the exogenous next state and the endogenous current state and action conditioned on the exogenous current state.
+8.3. Positioning our work in the literature
+While previous studies like Efroni et al.
+(2022); Dietterich et al.
+(2018) focus on the decomposition of states into endogenous and exogenous components, our work addresses a fundamentally different problem: the impact of integrating exogenous event information into reinforcement learning.
+Changing dynamics of environments based on externally specified “contexts” has also been studied under Contextual MDPs (Hallak et al., 2015).
+However, these studies consider contexts, where each context is constant for each episode and has a finite number of possible values.
+The aforementioned work proposes the CECE algorithm that learns a set of policies corresponding to all the contexts and determines the latent context for an episode (if unknown) to choose the policy corresponding to the current context.
+More specifically, CECE clusters an initial set of trajectories into groups corresponding to a fixed set of latent contexts.
+For each new episode, a partial realization is used to classify it into one of these clusters and to use a model learned on that cluster to exploit the rest of the trajectory.
+This algorithm cannot handle the setting considered in this paper because each new time step results in a change in the context, and knowledge of a latent context for the current time step does not guarantee the same context for the rest of the episode, except in a trivial special case.
+Modi et al.
+(2018) consider a relatively closer setting to ours where the context is known and possibly continuous, with their algorithm being more amenable to being adapted to handle dynamic contexts.
+However, it works by keeping count of the number of occurrences of $(s,a)$ in each of a set of representative contexts to learn a set of explicit models, which is only possible in the setting of finite state and action spaces.
+Tennenholtz et al.
+(2023) extend Contextual MDPs to handle dynamic contexts that are known to the agent, change at each time step, and influence the state transition distribution.
+However, this work differs from ours in several significant aspects.
+This work considers the contexts at each time step to depend on the history of states, actions, and contexts, while the next state depends only on the current state, action, and context.
+This means causality exists in both directions, from states to contexts and vice versa.
+Whereas in our work, the contexts are exogenous events that cannot be controlled by the agent, and the causality flows only in one direction, from contexts to states.
+Further, this work considers finite state spaces, and their analysis is explicitly geared towards such spaces. Their bounds depend on the size of the state and action sets.
+In Tennenholtz et al.
+(2023), the context space is a fixed finite set of $M$ vectors, and a latent map is learned from the (state, action, context) space to $\mathbb{R}^{M}$, with the range of the map still being a finite albeit exponentially increasing set.
+This work also considers a finite horizon MDP.
+This relatively simple setting allows one to tractably find the latent map and learn an optimal policy.
+In this sense, our setting is much more general and challenging.
+In addition to this, the work of Tennenholtz et al.
+(2023) presents regret analysis, while our study focuses on the convergence results of algorithms.
+Our analysis addresses the non-stationarity induced in the MDP due to the influence of non-Markovian events, necessitating the learning of policies that are contingent upon both the current state and a history of prior events.
+Although this approach bears resemblance to the strategies employed in Partially Observable MDPs (POMDPs) (Sunberg and Kochenderfer, 2018), the underlying rationale diverges significantly.
+In POMDPs, histories of observations are utilized despite the Markovian nature of state transition and observation functions, primarily to mitigate the agent’s inability to perceive the true state.
+Conversely, in our framework, both the state and external events are fully observable, and the necessity for an augmented space that incorporates event histories arises from the intrinsic non-Markovian characteristics of the non-stationary environment.
+Moreover, the causal framework of our scenario is distinctly different.
+The events exhibit non-Markovian properties, and the state at any given time directly affects the distribution of the subsequent state, regardless of the knowledge of the event.
+In POMDPs, while beliefs and sufficient statistics are contingent upon a history of observations, the states themselves maintain a Markovian nature, leading to observations that are causally independent.
+Our paper makes significant inroads into understanding how characteristics of exogenous event processes affect the tractability of the problem and the convergence and sample complexity of RL algorithms.
+To the best of our knowledge, there are no existing works in the literature that offer these findings.
 9. Discussion and Future Directions
 State space
-The theorem proposed is general and relies on learning functions in an augmented state space that blows up in the presence of long-term dependencies due to the external events.
-For a practical algorithm, it is necessary to compress this augmented state space to something more manageable.
-This can be done implicitly by using function approximators such as recurrent neural networks or LSTMs, or explicitly by learning compressed state representations using mutual information bottlenecks etc.
-While these are practical methods, it is also an open problem on how to theoretically analyze such a learnt compressed state representation.
+This work sets the stage for understanding problems related to RL under the influence of external temporal events and provides several theoretical insights.
+The results are established in a very general setting and rely on learning functions in an augmented state space that is exponentially bigger in the presence of long-term dependencies due to external events.
+To mitigate this, one needs to develop provable approximate methods by reducing the size of the state space.
+Due to the ”factorization” of the environment into Markovian states and non-Markovian events that impact these states, compressing the histories of events into latent features that are independent of the states may provide a strategic advantage.
+This approach could facilitate the learning of representations of the expanded states by developing a distinct event distribution model that can be scaled up, as it does not rely on the agent, alongside a separate state transition model that necessitates agent interaction with the environment.
+A compressed representation of events, derived from extensive event data, can subsequently be utilized to learn a feasible state transition kernel.
+This can be achieved either implicitly through function approximators such as recurrent neural networks (or LSTMs), or explicitly by learning compressed state representations using mutual information bottlenecks, etc.
+While these are practical methods, the theoretical analysis of such learned compressed state representations remains an open problem.
 Generalization
-The sample complexity analysis in Section 7 hinges on generalizing sample error of policy evaluation to an expected error over the entire state space.
+The analysis of sample complexity in Section 6 hinges on generalizing sample error of policy evaluation to an expected error over the entire state space.
 Such generalization generally requires assumptions on the distribution over states.
-In the theorem stated by us, the assumption was on the rate of convergence of the Markov chain induced from the augmented MDP by the policy under evaluation, which is common in the reinforcement learning literature.
-However, in our case, the specific structure of the state space, i.e, separation of the augmented state space into states and events with different characteristics, may lead itself to a different types of assumptions.
-For example, the recurrence and aperiodicity of the Markov chain may not be a good assumption, but a stationarity assumption on just the event sequence might be justified.
+In our result, the assumption pertains to the rate of convergence of the Markov chain induced from the augmented MDP by the policy under evaluation, a notion frequently encountered in reinforcement learning literature.
+However, in our case, the specific structure of the state space–specifically, the separation of the augmented state space into states and events with different characteristics–may necessitate alternate assumptions.
+For instance, while the recurrence and aperiodicity of the Markov chain may not hold, it may be reasonable to assume stationarity solely for the event sequence.
 Due to this, the generalizability of the value function in our setting can be analyzed in a more specific way that takes into account this exogenous structure.
@@
 The policy improvement step involves optimizing over the action space based on the expected one-step reward and the value function at the next time step, which requires knowledge of the model.
-Thus, in the case of an unknown transition and reward model, it is important to investigate the sample complexity of learning the models from samples.
-Appendix A Proofs of lemmas
-A.1. Proof of Lemma  5.2
-We adapt the proof of Theorem 1 from Xiao et al. (2019).
-We wish to study the value function of the policy in two different MDPs $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$.
-For convenience, without any loss in generality, we can extend the state space of $\mathcal{M}^{(T)}$ to that of $\mathcal{M}_{X}$ by including the information of every past event in the state, without changing the transition function.
-We define another class of auxiliary MDPs $\mathcal{M}^{(T)}_{h}$, $h\in 0\cup\mathbb{N}$, in which when starting from initial (augmented) state $\bar{s}\in\bar{S}$, the transitions follow the transition kernel of $\mathcal{M}_{X}$ for $h$ time steps and then the transition kernel of $\mathcal{M}^{(T)}$ for all transitions thereafter.
-The value functions induced by $\pi$ in $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}$ can be related using these new intermediate MDPs, which interpolate between those two environments.
-Since only one policy is under discussion, the $\pi$ in $V^{\pi}$ is omitted for the rest of this proof for simplicity of notation.
-$V$ denotes the value function of the policy in $\mathcal{M}_{X}$, $V^{(T)}$ denotes its value in $\mathcal{M}^{(T)}$, and $V_{h}^{(T)}$ in $\mathcal{M}^{(T)}_{h}$.
-The following relations follow from the definitions of the auxiliary and intermediate MDPs and the boundedness of $r$:
+Thus, in the case of an unknown transition and reward model, it is important to investigate the sample complexity of learning the models from samples and how this affects the performance of the resultant policy.
+Appendix A Notation
+Table 1. Table of notation
+$\mathcal{M}$ Stationary MDP
+$S$ State space
+$A$ Action space
+$Q$ Transition probability kernel
+$r$ Reward function
+$r(s,a)$ Reward for taking action $a$ at state $s$
+$r(s,a,s^{\prime})$ Reward if taking action $a$ at state $s$ led to next state $s^{\prime}$
+$\gamma$ Discount factor
+$\pi$ Policy
+$V^{\pi}$ Value function of policy $\pi$
+$\mathcal{T}$ Optimal Bellman operator
+$\pi^{*}$ Optimal policy
+$V^{*}$ Value function of $\pi^{*}$
+$\mathsf{TV}(.\;,\;.)$ Total Variation distance between two probability distributions
+$\mathcal{X}$ Space of event marks
+$\mathcal{M}_{X}$ Augmented MDP
+$\mathcal{M}^{(T)}_{X}$ Approximate augmented MDP with finite history of $T+1$ events
+$H$ History of external events till time $t$
+$Q_{H}$ Transition kernel of the MDP due to influence of event history $H$
+$Q^{X}_{H}$ Distribution of the mark of the event process due to history $H$
+$\bar{S}$ Augmented state space
+$\bar{s}$ State in $\bar{S}$
+$x$ Event mark, an element of $\mathcal{X}$
+$x_{0:t}$ Events that occurred withing the past $t$ time steps
+$x_{t:\infty}$ Events older than $t$ time steps
+$\left(x_{0:t},x^{\prime}_{t+1:\infty}\right)$ Event history obtained by concatenating past few events of $x$
+with older events from $x^{\prime}$
+$\varphi_{i}$ Basis functions
+$\phi$ Feature function consisting of all basis functions
+$\Phi$ Matrix containing feature vectors
+$d$ Dimensionality of feature representation, i.e., number of basis functions
+$\hat{P}$ Empirical transition matrix
+$\hat{V}$ Value function estimated using pathwise LSTD
+$\widetilde{V}$ $\hat{V}$ clipped at $\frac{1}{1-\gamma}$
+$\nu$ Lower bound on the eigenvalues of $\frac{1}{N}\Phi^{T}\Phi$
+Appendix B Proofs of lemmas
+B.1. Proof of Lemma  4.2
+The aim is to study the value function of the policy in two MDPs $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$, where $\mathcal{M}_{X}$ is the non-stationary MDP with augmented states consisting of the actual state and the entire history of events till then, and $\mathcal{M}^{(T)}_{X}$ is an approximate MDP where only events in the past $T$ time steps affect the state transition and event distribution.
+For convenience, without any loss in generality, one can extend the state space of $\mathcal{M}^{(T)}_{X}$ to that of $\mathcal{M}_{X}$ by including the information of every past event in the state, without changing the transition function.
+Now, We define a class of auxiliary MDPs $\mathcal{M}^{(T)}_{h}$, $h\in\left\{0,1,2,\dots\right\}$, in which when starting from initial (augmented) state $\bar{s}\in\bar{S}$, the transitions follow the transition kernel of $\mathcal{M}_{X}$ for $h$ time steps and then the transition kernel of $\mathcal{M}^{(T)}_{X}$ for all transitions thereafter.
+The value functions induced by $\pi$ in $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$ can be related using these new intermediate MDPs, which interpolate between those two environments.
+Since only a single policy is under discussion, the $\pi$ in $V^{\pi}$ is omitted for the rest of this proof for simplicity of notation.
+$V$ denotes the value function of the policy in $\mathcal{M}_{X}$, $V^{(T)}$ denotes its value in $\mathcal{M}^{(T)}_{X}$, and $V_{h}^{(T)}$ in $\mathcal{M}^{(T)}_{h}$.
+The following relations follow from the definitions of the auxiliary and intermediate MDPs and the boundedness of $r$.
 $V^{(T)}=V_{0}^{(T)},\quad V=\lim_{h\rightarrow\infty}V^{(T)}_{h}$
-and so,
+and hence,
 $V-V^{(T)}$ $=\lim_{H\rightarrow\infty}V_{H}^{(T)}-V_{0}$
@@
 $=\sum_{h=0}^{\infty}\left(V_{h+1}^{(T)}-V_{h}^{(T)}\right).$
-That is, the entire difference can be characterized in terms of differences in $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ for each $h$. Now for any $s\in\bar{S}$,
-$V_{h}^{(T)}(s)$ $=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]$
-$\quad+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
-$\quad+\sum_{t=h+1}^{\infty}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t-h}^{(T)}\circ p_{h}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p^{(T)}(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right],$
-where $p_{t}(.|s)$ denotes a transition of $t$ steps in the environment $\mathcal{M}_{X}$ starting from $s$, a superscript (T) denotes the corresponding object in $\mathcal{M}^{(T)}$, and the policy $\pi$ is implicit whenever it is not explicitly mentioned in any expression.
+That is, the difference between $V$ and $V^{(T)}$ can be characterized in terms of differences in $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ for each $h\in\{0,1,2,\dots\}$.
+Now, for any augmented state $\bar{s}\in\bar{S}$,
+$V_{h}^{(T)}(\bar{s})$ $=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum_{t=h+1}^{\infty}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t-h}^{(T)}\circ p_{h}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right],$
+where $p_{t}(.|\bar{s})$ denotes a transition of $t$ steps in the environment $\mathcal{M}_{X}$ starting from $\bar{s}$, a superscript (T) denotes the corresponding object in $\mathcal{M}^{(T)}_{X}$, and the policy $\pi$ is implicit whenever it is not explicitly mentioned in any expression.
+The reward function does not depend on the external events, so $r(\bar{s}_{t},a_{t},\bar{s}_{t+1)}$ is just $r(s_{t},a_{t},s_{t+1})$, where $\bar{s}=(s,x)$ and $s$ is the actual state.
 This can also be written as
-$V_{h}^{(T)}(s)=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}s_{t}\sim p_{t}(.|s)\\ a_{t}\sim\pi(.|s_{t})\\ s_{t+1}\sim p(.|s_{t},a_{t})\end{subarray}}\left[r(s_{t},a_{t},s_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h}).$
+$V_{h}^{(T)}(\bar{s})=\sum_{t=0}^{h-1}\gamma^{t}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{t}\sim p_{t}(.|\bar{s})\\ a_{t}\sim\pi(.|\bar{s}_{t})\\ \bar{s}_{t+1}\sim p(.|\bar{s}_{t},a_{t})\end{subarray}}\left[r(\bar{s}_{t},a_{t},\bar{s}_{t+1})\right]+\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h}).$
 Using these two representations for $V_{h}^{(T)}$ and $V_{h+1}^{(T)}$ respectively gives
-$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{h+1}\sim p(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
-$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p_{h+1}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
-$-\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\\ s_{t+1}\sim p^{(T)}(.|s_{h},a_{h})\end{subarray}}\left[r(s_{h},a_{h},s_{h+1})\right]$
-$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h+1}\sim p^{(T)}\circ p_{h}(.|s)\end{subarray}}V^{(T)}(s_{h+1})$
-$=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})$
-$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}r(s_{h},a_{h},s_{h+1})\bigg{]}$
-$+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}s_{h}\sim p_{h}(.|s)\\ a_{h}\sim\pi(.|s_{h},a_{h})\end{subarray}}\bigg{[}\mathsf{E}_{s_{h+1}\sim p(.|s_{h},a_{h})}V^{(T)}(s_{h+1})$
-$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{s_{h+1}\sim p^{(T)}(.|s_{h},a_{h})}V^{(T)}(s_{h+1})\bigg{]}.$
-In the above equations, $s,s_{t}$ etc., denote the augmented state, which consists of the state and event mark history till then.
+$V_{h+1}^{(T)}(\bar{s})-V_{h}^{(T)}(\bar{s})$ $=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
+$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h+1}\sim p_{h+1}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h+1})$
+$\qquad\qquad-\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\\ \bar{s}_{t+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})\end{subarray}}\left[r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\right]$
+$\qquad\qquad\qquad\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h+1}\sim p^{(T)}\circ p_{h}(.|\bar{s})\end{subarray}}V^{(T)}(\bar{s}_{h+1})$
+$=\gamma^{h}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h})\end{subarray}}\bigg{[}\mathsf{E}_{\bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})}r(\bar{s}_{h},a_{h},\bar{s}_{h+1})$
+$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{\bar{s}_{h+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})}r(\bar{s}_{h},a_{h},\bar{s}_{h+1})\bigg{]}$
+$\qquad\qquad+\gamma^{h+1}\mathsf{E}_{\begin{subarray}{c}\bar{s}_{h}\sim p_{h}(.|\bar{s})\\ a_{h}\sim\pi(.|\bar{s}_{h},a_{h})\end{subarray}}\bigg{[}\mathsf{E}_{\bar{s}_{h+1}\sim p(.|\bar{s}_{h},a_{h})}V^{(T)}(\bar{s}_{h+1})$
+$\qquad\qquad\qquad\qquad\qquad-\mathsf{E}_{\bar{s}_{h+1}\sim p^{(T)}(.|\bar{s}_{h},a_{h})}V^{(T)}(\bar{s}_{h+1})\bigg{]}.$
 Since the current state and the current event mark are independent and depend only on the previous events, the transition kernel can be factored as two independent distributions over $S$ and $\mathcal{X}$.
-So, replacing the augmented state $s\in\bar{S}$ with the explicit representation $(s,H)$, where $s\in S$ and $H\in\mathcal{X}^{\infty}$ gives, for any bounded measurable function $f$ on $\mathbb{K}\times\mathcal{X}^{\infty}$,
+Hence, replacing the augmented state $\bar{s}\in\bar{S}$ with the explicit representation $(s,H)$, where $s\in S$ and $H\in\mathcal{X}^{\infty}$ gives, for any bounded measurable function $f$ on $S\times\mathcal{X}^{\infty}$,
 $\mathsf{E}_{(s_{t+1},H_{t+1})\sim p(.|s_{t},H_{t},a_{t})}f(s_{t+1},H_{t+1})-\mathsf{E}_{(s_{t+1},H_{t+1})\sim p^{(T)}(.|s_{t},H_{t},a_{t})}f(s_{t+1},H_{t+1})$
-$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
-$=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
-$+\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
-$\leq\mathsf{E}_{\begin{subarray}{c}X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}\left[\left|\left|f(.,X_{t+1},H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q_{H_{t}}(.|s_{t},a_{t}),Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
-$+\mathsf{E}_{s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})}\left[\left|\left|f(s_{t+1},.,H_{t})\right|\right|_{\infty}\mathsf{TV}\left(Q^{X}_{H_{t}}(.|s_{t},a_{t}),Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
-$\leq||f||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+\sum_{t=T+1}^{\infty}N_{t}\right).$
-The second to last inequality is a result of Total Variation Distance being an integral probability metric (Sriperumbudur et al., 2009). Using this expression in the previous equation gives
-$V_{h+1}^{(T)}(s)-V_{h}^{(T)}(s)$ $\leq\gamma^{h}||r||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
-$\qquad\qquad+\gamma^{h+1}||V^{(T)}||_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
-$\leq\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
+$\qquad=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
+$\qquad=\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
+$\qquad\qquad\qquad+\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})-\mathsf{E}_{\begin{subarray}{c}s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\\ X_{t+1}\sim Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}f(s_{t+1},X_{t+1},H_{t})$
+$\qquad\leq\mathsf{E}_{\begin{subarray}{c}X_{t+1}\sim Q^{X}_{H_{t}}(.|s_{t},a_{t})\end{subarray}}\left[\left\|f(.,X_{t+1},H_{t})\right\|_{\infty}\mathsf{TV}\left(Q_{H_{t}}(.|s_{t},a_{t}),Q^{(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
+$\qquad\qquad\qquad+\mathsf{E}_{s_{t+1}\sim Q^{(T)}_{H_{t}}(.|s_{t},a_{t})}\left[\left\|f(s_{t+1},.,H_{t})\right\|_{\infty}\mathsf{TV}\left(Q^{X}_{H_{t}}(.|s_{t},a_{t}),Q^{X(T)}_{H_{t}}(.|s_{t},a_{t})\right)\right]$
+$\qquad\leq\left\|f\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+\sum_{t=T+1}^{\infty}N_{t}\right).$
+The second to last inequality is a result of Total Variation Distance being an integral probability metric (Sriperumbudur et al., 2012). Using this expression in the previous equation gives
+$V_{h+1}^{(T)}(\bar{s})-V_{h}^{(T)}(\bar{s})$ $\leq\gamma^{h}\left\|r\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
+$\qquad\qquad+\gamma^{h+1}\left\|V^{(T)}\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
+$\leq\gamma^{h}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
 This gives the final result,
-$||V-V^{(T)}||_{\infty}$ $\leq\sum_{h=1}^{\infty}\gamma^{h}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
-$=\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{(T)}||_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
-A.2. Proof of Lemma 6.1
+$\left\|V-V^{(T)}\right\|_{\infty}$ $\leq\sum_{h=1}^{\infty}\gamma^{h}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
+$=\frac{1}{1-\gamma}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{(T)}\right\|_{\infty}\right)\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right).$
+B.2. Proof of Lemma 5.1
+Let $\pi$ be a policy that depends only on the current state and the past $T$ events.
 $V^{\pi}((s,x_{0:\infty}))$ $=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:\infty}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:\infty}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left[r(s,a,s^{\prime})+V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))\right],$
 $V^{\pi}((s,x_{0:T},0))$ $=\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left[r(s,a,s^{\prime})+V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right].$
-Taking the difference between these two expressions, and adding and subtracting the additional term
+Taking the difference between these two expressions and adding and subtracting the additional term
 $\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))$
@@
 $\left|V^{\pi}((s,x_{0:\infty}))-V^{\pi}((s,x_{0:T},0))\right|$
-$\leq||r||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}\right)+||V^{\pi}||_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+N_{t}\right)$
-$\qquad\qquad+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left|V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))-V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right|.$
+$\qquad\qquad\qquad\leq\left\|r\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}\right)+\left\|V^{\pi}\right\|_{\infty}\left(\sum_{t=T+1}^{\infty}M_{t}+N_{t}\right)$
+$\qquad\qquad\qquad\qquad\qquad+\gamma\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T}}(.|s,a)\\ x^{\prime}\sim Q_{x_{0:T}}(.)\\ a=\pi((s,x_{0:T}))\end{subarray}}\left|V^{\pi}((s^{\prime},x^{\prime},x_{0:\infty}))-V^{\pi}((s^{\prime},x^{\prime},x_{0:T}))\right|.$
 In the above inequality, the left-hand side is the difference between two value functions whose first $T$ events are the same and differ everywhere else, but the right-hand side has a discounted term in which the first $T+1$ terms coincide.
 So, repeating the same procedure results in value function differences at states that increasingly coincide.
-Since each repetition adds a $\gamma$ factor to the term, which itself is bounded by $\frac{1}{1-\gamma}$ because of bounded rewards, the resulting series converges, giving us the following bound:
+Since each repetition adds a $\gamma$ factor to the term, which itself is bounded by $\frac{1}{1-\gamma}$ because of bounded rewards, the resulting series converges, giving us the following bound.
 $\sup_{\bar{s}=(s,x_{0:\infty})}\left|V^{\pi}((s,x_{0:\infty}))-V^{\pi}((s,x_{0:T},0))\right|$
-$\leq\sum_{t=T+1}^{\infty}\gamma^{t-T-1}\left[||r||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}\right)+||V||_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}+N_{t^{\prime}}\right)\right]$
-$\leq\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}\sum_{t^{\prime}=t}^{\infty}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
-$\leq\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\sum_{t=0}^{T-t^{\prime}-1}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
-$=\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\frac{1-\gamma^{t^{\prime}-T}}{1-\gamma}(M_{t^{\prime}}+N_{t^{\prime}})\leq\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
-Appendix B Details of Theorem 7.1
+$\qquad\qquad\qquad\qquad\leq\sum_{t=T+1}^{\infty}\gamma^{t-T-1}\left[\left\|r\right\|_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}\right)+\left\|V\right\|_{\infty}\left(\sum_{t^{\prime}=t}^{\infty}M_{t^{\prime}}+N_{t^{\prime}}\right)\right]$
+$\qquad\qquad\qquad\qquad\leq\left(1+\frac{\gamma}{1-\gamma}\right)\sum_{t=T+1}^{\infty}\sum_{t^{\prime}=t}^{\infty}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
+$\qquad\qquad\qquad\qquad\leq\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\sum_{t=0}^{T-t^{\prime}-1}\gamma^{t}(M_{t^{\prime}}+N_{t^{\prime}})$
+$\qquad\qquad\qquad\qquad=\frac{1}{1-\gamma}\sum_{t^{\prime}=T+1}^{\infty}\frac{1-\gamma^{t^{\prime}-T}}{1-\gamma}(M_{t^{\prime}}+N_{t^{\prime}})$
+$\qquad\qquad\qquad\qquad\leq\frac{1}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
+Appendix C Details of Theorem 6.1
 Intuitively, bounding the expected error of the estimated value function based on the error at just a few samples requires the samples to be close to the stationary distribution, and the function values at the samples to be close to their expected values.
@@
 Definition 1.
-A Markov chain $\mathcal{M}=\left(X_{t}\right)_{t\geq 1}$ is said to be exponentially $\beta$-mixing with parameters $\bar{\beta},b,\kappa$ if its mixing coefficients
+A Markov chain $\left(X_{t}\right)_{t\geq 1}$ is said to be exponentially $\beta$-mixing with parameters $\bar{\beta},b,\kappa$ if its mixing coefficients
 $\beta_{i}=\sup_{t\geq 1}\left[\sup_{B\in\sigma\left(X_{t+i},\dots\right)}\left|\mathsf{P}\left(B|X_{1},\dots,X_{t}\right)-\mathsf{P}(B)\right|\right]$
 satisfy $\beta_{i}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right)$.
-When such an exponentially mixing Markov chain is also ergodic and and aperiodic with stationary distribution $\rho$, for any initial distribution $\lambda$, there is a bound on the following total variation: (Definition 21 of Lazaric et al.
+When such an exponentially mixing Markov chain is also ergodic and and aperiodic with stationary distribution $\rho$, for any initial distribution $\lambda$, there is a bound on the following total variation: (Definition 21 in Lazaric et al.
 (2012))
-$\left|\left|\int_{\mathcal{X}}\lambda(dx)\mathsf{P}\left(.|x\right)-\rho(.)\right|\right|_{\mathsf{TV}}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right).$
+$\left\|\int_{\mathcal{X}}\lambda(dx)\mathsf{P}\left(.|x\right)-\rho(.)\right\|_{\mathsf{TV}}\leq\bar{\beta}\exp\left(-bi^{\kappa}\right).$
 This helps in bounding the difference between the norm of a function in the stationary distribution and the empirical distribution defined using a few samples.
-Lemma B.1 (Generalization Lemma for Markov chains, Lemma 24 of Lazaric et al. (2012)).
-Let $\tilde{\mathcal{F}}$ be the $d$-dimensional class of linear functions truncated at threshold $B$.
+Lemma C.1 (Generalization Lemma for Markov chains, Lemma 24 of Lazaric et al. (2012)).
+Let $\widetilde{\mathcal{F}}$ be a $d$-dimensional class of linear functions truncated at threshold $B$.
 Let $\left(X_{t}\right)_{t\in[n]}$ be a sequence of samples from an ergodic, aperiodic exponentially $\beta$-mixing Markov chain with parameters $\bar{\beta},b,\kappa$, arbitrary initial distribution, and stationary distribution $\rho$.
-If the first $\tilde{n}$ samples are discarded and only the last $n-\tilde{n}$ samples are used, then with probability at least $1-\delta$, the empirical and $l_{2}(\rho)$ norm of the any function $\tilde{f}\in\tilde{\mathcal{F}}$ are related as,
-$||\tilde{f}||-2||\tilde{f}||_{X_{1:n}}$ $\leq\epsilon(\delta),$
-$||\tilde{f}||_{X_{1:n}}-2\sqrt{2}||\tilde{f}||$ $\leq\epsilon(\delta),$
-$\text{for }\epsilon(\delta)=12B\sqrt{\frac{2\Lambda(n-\tilde{n},d,\delta)}{n-\tilde{n}}\max\left\{\frac{\Lambda(n-\tilde{n},d,\delta)}{1},1\right\}^{\frac{1}{\kappa}}},$
+If the first $\widetilde{n}$ samples are discarded and only the last $n-\widetilde{n}$ samples are used, then with probability at least $1-\delta$, the empirical and $l_{2}(\rho)$ norm of the any function $\widetilde{f}\in\widetilde{\mathcal{F}}$ are related as,
+$\|\widetilde{f}\|-2\|\widetilde{f}\|_{X_{1:n}}$ $\leq\epsilon(\delta),$
+$\|\widetilde{f}\|_{X_{1:n}}-2\sqrt{2}\|\widetilde{f}\|$ $\leq\epsilon(\delta),$
+$\text{for }\epsilon(\delta)=12B\sqrt{\frac{2\Lambda(n-\widetilde{n},d,\delta)}{n-\widetilde{n}}\max\left\{\frac{\Lambda(n-\widetilde{n},d,\delta)}{1},1\right\}^{\frac{1}{\kappa}}},$
 $\Lambda(n,d,\delta)=2(d+1)\log n+\log\frac{\epsilon}{\delta}+\log^{+}\left(\max\left\{16(6e)^{2(d+1)},\bar{\beta}\right\}\right),$
@@
 (2012) derived the following probabilistic lower bound on the smallest eigenvalue of the sample-based Gram matrix as a function of the smallest eigenvalue of the Gram matrix $G=\mathsf{E}_{x\sim\rho}\left[\phi(x)\phi(x)^{T}\right]$.
-Lemma B.2 (Lemma 4 of Lazaric et al. (2012)).
+Lemma C.2 (Lemma 4 in Lazaric et al. (2012)).
 Let $\omega>0$ be the smallest eigenvalue of the Gram matrix defined above.
@@
 $\Lambda(n,d,\delta)=2(d+1)\log n+\log\frac{e}{\delta}+\log^{+}\left(\max\left\{18(6e)^{2(d+1)},\bar{\beta}\right\}\right)$
-comes from a simpler version of Lemma B.1 without the additional initial $\tilde{n}$ samples for burn-in of the Markov chain.
-The generalization Lemma B.1, combined with the lower bound on the eigenvalues and our results in Section 6, give rise to the proof of the following result on the sample complexity of policy evaluation, detailed in Section 7.
-B.1. Theorem 7.1
+comes from a simpler version of Lemma C.1 without the additional initial $\widetilde{n}$ samples for burn-in of the Markov chain.
+The generalization Lemma C.1, combined with the lower bound on the eigenvalues and our results in Section 5, give rise to the proof of the following result on the sample complexity of policy evaluation, detailed in Section 6.
+C.1. Theorem 6.1
 Assume that the policy $\pi$ induces from the MDP $\mathcal{M}_{X}$ a $\beta$-mixing Markov chain with parameters $\beta$, $b$, $\kappa$.
-Let $\bar{s}_{1},\dots,\bar{s}_{N+\tilde{N}}$ be a sample path obtained using the policy $\pi$.
-Suppose the first $\tilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\tilde{V}$ of the true value function $V$.
-Then, provided the number of discarded samples is $\tilde{N}=\left(\frac{1}{b}\log\frac{2e\bar{\beta}n}{\delta}\right)^{\frac{1}{\kappa}}$, with probability at least $1-\delta$,
-$||\tilde{V}-V||_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+||\Pi^{trunc}V-\Pi V||_{\rho}\right)$
+Let $\bar{s}_{1},\dots,\bar{s}_{N+\widetilde{N}}$ be a sample path obtained using the policy $\pi$.
+Suppose the first $\widetilde{N}$ samples are discarded for Markov chain burn-in, and the remaining $N$ are used to compute the truncated least-squares path-wise estimate $\widetilde{V}$ of the true value function $V$.
+Then, provided the number of discarded samples is $\widetilde{N}=\left(\frac{1}{b}\log\frac{2e\bar{\beta}n}{\delta}\right)^{\frac{1}{\kappa}}$, with probability at least $1-\delta$,
+$\left\|\widetilde{V}-V\right\|_{\rho}$ $\leq\frac{4\sqrt{2}}{\sqrt{1-\gamma^{2}}}\left(\frac{3}{(1-\gamma^{2})}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)+\left\|\Pi^{trunc}V-\Pi V\right\|_{\rho}\right)$
 $\qquad\qquad+\frac{2L}{(1-\gamma)^{2}}\sqrt{\frac{d}{\nu}}\left(\sqrt{\frac{2\log(2d/\delta)}{N}}+\frac{1}{N}\right)$
@@
 $\epsilon_{2}(N,\delta)$ $=12\left(\frac{1}{1-\gamma}+L||\alpha^{*}||\right)\sqrt{\frac{2\Lambda_{2}(N,\delta/4)}{N}\max\left\{\frac{\Lambda_{2}(N,\delta/4)}{b},1\right\}^{\frac{1}{\kappa}}},$
-with $\Lambda_{1}(N,d,\delta)$ and $\Lambda_{2}(N,\delta)$ as defined in Lemma B.1, and $\alpha^{*}$ such that $\Pi V=\sum_{i}\alpha^{*}_{i}\varphi_{i}$.
+with $\Lambda_{1}(N,d,\delta)$ and $\Lambda_{2}(N,\delta)$ as defined in Lemma C.1, and $\alpha^{*}$ such that $\Pi V=\sum_{i}\alpha^{*}_{i}\varphi_{i}$.